A Bibliography of Publications in Notices of the American Mathematical Society (1990–1999)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 August 2024 Version 1.00

Title word cross-reference

$ [KR97]. J(τ) [Kno90]. O [Knu98]. p [Tei95].

-adic [Tei95].

01015b [Var99].

1990 [Ano90a, Ano90m, Ano90b, Ano90c]. 1990s [Ano90l, TW99]. 1991 [Ano91a, Ano91b, Ano91c]. 1992 [Ano92a, Ano93a]. 1993 [Ano93b, Ano93c]. 1994 [Ano94c, Ano94a]. 1995 [Ano95a, Ano95b, Ano95c, Ano95d]. 1996 [Ano96a, Ano96b, Dav97b]. 1997 [Ano97a, Ano97c, Ano97b, DMR98]. 1998 [LLMM99]. 1999 [Ano99a, Ano99c, Ano99b].

20th [Lan95c]. 21st [Ril98]. 269 [Kra94].

3.0 [Dar93, Har93].

41 [Bae94, Hor94]. 42 [Seg96].

90 [GJM^{+}90]. 90s [Oke96]. 90th [Duf97]. 93j [Var99]. 94 [Jac94b]. 95d [Kra94]. 95j [JM97]. 95m [Far96a]. 96c [Rad96]. 96f [Seg96]. 98f [Dub99]. 99g [Jer99]. 99j [Mil98].

A. [CDK^{+}95, Fal95]. academia [HJT98, Mac98]. academic [Bil91, JM97]. accomplishments [Ano90l]. according [Jer99]. acrobatics [Jac97a]. action [Sum95]. Addendum [AV90]. address [Mor99]. adiabatic [Uhl95]. adic [Tei95]. Advances [Grä93b, Grä94c, Grä94a, Grä94b]. adventures [Mac97b]. Africa [LS97]. after [Hel97, Mac98]. Age [Hwa95]. agency [Sha92]. ago [Ewi96]. Ahlfors [BEH^{+}98, GKO98, Oss99]. Aided [BB92]. Alain [JM97]. Alan [Ano90d]. Albert [GHS97]. Alberto [Ano92b, CKSW98]. Algebra [Eck99, Sto91, Ano92a, Lan97b, Ano90m, Ano95a, Mac97d, Mac97c]. algebraic [Bor99, Ray99]. algebras [Rei97]. algorithms [Lan95a, Lan95b]. Alice [Ano95c]. Allyn [RA96]. Almgren [Whi97, Mac97b]. alternating [BP99]. Ambrose [SW96]. Amer [Bae94, Hor94, Kra94, Seg96]. American [Jac99a]. AMS [Dav97b, DMR98, Ano90k, Ano98a, How98, Jac94a, Jac97d, Jac99d]. Analysis [BB92, Eck99, Str99, Jac95, Kna96a, Kna96b, Wal97]. André [BBCB^{+}99, BCC^{+}99, Bor99, Car99, Kna99, Ray99, Shi99, Var99]. Andreas [HWZ91]. Andrew [Coa96]. announced [Dar99]. Annual [DMR98, Dav97b, DMR99]. Antoni [CS92]. anyway [Oke96]. Aperiodic [Kup99]. Appendix [Ano90l]. applications [Dau95]. Applied [Ano95b, You97a, Ano90a]. apprenticeship [Var99]. approach [Imm95, McD98]. Arithmetic [Tur91]. Arnol’d [Lui97]. Arthur [Jac97d, Jac99d]. article [Dev92a, Oke96]. aspirations [LS97]. Association [Blu91, TW99]. Attacks [Bon99]. Auslander [CKK^{+}98, PR95]. Author [Hor94]. autobiography [Cha98]. Automatic [Ano91d]. aux [Cha98]. award [Coa96, Ano93g]. awarded [Ano90a, Ano90m, Ano90b, Ano91e, Ano91a, Ano91b, Ano91c, Ano92c, Ano95a, Ano95b, Ano95c, Ano95d, Ano99b]. awards [Ano91d, Ano91j]. AWM [TW99].

B [Jer99, Jac99e, Ano90l]. B. [Jer99]. Bôcher [Ano94c, Ano99a]. Bargmann [LSW90]. Barry [KS98]. Bartel [DS97]. Basel [Var99]. battles [Jac97b, Jac97c]. BBC [Gra97]. beacon [Dol95]. Beal [Mau97]. beautiful [Mil98]. Beauty [Pic91]. been [Rot97b]. Being [Rei96]. Bergman [Ano91e, Ano92c, Ano93h]. Bernadette [Ano99b]. Bernard [KT99]. Berndt [Ano96b]. Bers [MAC^{+}95b]. between [Hag96]. beyond [Oss99]. Big [Knu98]. biographical [RA96]. biology [Hop95]. Birkhäuser [Var99]. Birkhoff [Lan97b, Mac97c, You97a]. birthday [Duf97]. Bitmap [Cas93]. Bitmap-Free [Cas93]. Black [Chi96]. Boas [Fis92]. Bonn [Jac98a]. Book [Rad96]. Borcherds [Jac98b]. Boris [Ano93d]. both [Run97]. Bourbaki [Bor98]. Branges [Ano90f]. Bremermann [RA96]. brewery [Jac97a]. brief [Blu91]. Bruce [Ano96b]. budget [Ano98b, Tho97]. building [Jac98a, Ril98].

C [Ano96b, Dev92a]. CA [JM97]. Calculating [Leh94]. Calculus [KR97, Knu98, Mum97, Gil97]. Calderón [Ano92b, CKSW98, Ste98]. California [Jac97b, Jac97c]. Cambridge [Rad96]. came [Sha92]. Canada [Jac92]. Capabilities [Mil90]. Capillary [Fin99]. career [Mac98]. careers [Ano91h, BHL^{+}91]. Carl [CFV96]. Carleson [Ano92d]. Carroll [Pri90]. Cartan [Jac99e]. Cartwright [MT99]. causes [Jac90]. celebrates [Duf97]. celebration [Dau95]. century [ADR97, Bro97, Ewi96, Lan95c, Ril98]. Cesari [Ano90e, Ano90j]. Challenges [Bro97]. chance [RS99]. Chang [Ano95c]. change [Roi91]. Changes [Max97]. chaos [FG95]. chaotic [You98]. Charles [Ano92c]. chemistry [Hag96]. Chern [JK98]. Chih [DPR^{+}98]. Chih-Han [DPR^{+}98]. Chinese [Jac97a]. Chow [WCA^{+}96]. Chowla [AHW98]. Ciprian [Ano95b]. citizen [Col97]. classification [Sol95]. classroom [Pic91]. Clay [Jac99g]. cohomology [Ade97, Kno90]. cold [Sha92]. Cole [Ano90m, Ano92a, Ano95a, Ano97a, Kei94]. color [Tho98]. Columbus [Ano90a, Ano90b]. Column [DW92, CVG94]. combinatorial [ABT97]. coming [Gro98]. comment [Ano91h, BHL^{+}91, Hor94, Kra94]. commercial [Bes98]. comparative [CVG94]. complex [Eas99, Jac95, Lan97a]. complexity [Imm95, Saa95]. composition [Max97]. compression [Bar96]. comprise [CVG94]. computation [Imm95]. computational [Ser97]. Computer [BB92, Sto91]. Computers [Bar90, CHM94, Dev92b, DW92, Dev93, Dev94a, Dev94b, Dev94c, Dev94d, Dev94e]. conditions [Kob91]. Conference [Jac90]. conjecture [BP99, Dar99, Lan95d, Mau97, Rib93, Sar95]. Connes [JM97]. consciousness [Far96b]. contact [Ber98]. contemporary [Art99]. contexts [Par96]. Control [Grä94c]. convex [Goo98]. Copyright [Oke96]. Corrections [Seg96]. countries [LS97]. covering [YH97]. Coxeter [Duf97]. CPU [Doo90]. Crafoord [ST94]. Crimes [Sto91]. cross [Ni98]. cross-diffusion [Ni98]. Crossing [Hag96]. Crow [CVG94]. Crucial [Art99]. Cryptosystem [Bon99]. Culture [Zei93]. curtain [Sum95]. curves [Mir99].

D [AGS94, Ano93g]. Daniel [Ano96b, Asc92, Wil97]. database [Woo90]. David [Ano95a]. dead [Kot95]. Deane [Bor92]. death [Kra94, Hor94]. Deaths [Ano91f]. decades [Gro98]. DeGroot [Ber90]. demise [Jac98c]. departments [Ano97e, GY90]. Derrick [Bri93]. Descriptive [Imm95]. Developments [You98, Ade97]. Devlin [CHM94]. diagrams [Rei97]. Dickson [Kap93]. Diego [JM97]. difference [Jac90]. Differential [Mag99, Way97]. Diffusion [Ni98]. Diophantine [Lan95c]. Direction [Jac94a]. Dmitrii [Ano90g]. Do [OP97]. Dobrushin [MSV96]. doctoral [Ano92e]. doctorates [Jac91]. Does [AMS90]. dogmatic [Col97]. Donaldson [ST94]. doors [Jac92]. Doubling [Tor97]. dream [Jac99b, RS99]. Dwork [KT99]. dynamical [Kup99]. dynamics [Kir96, You98]. Dynkin [Rei97].

economics [Saa95]. Edgar [Ano90h]. Edition [DDW93]. Éditions [Cha98]. Editor [Knu98]. Educating [Ano92e]. education [Ano92g, Ano97d, Art99, Jac97b, Jac97c, Ril98, Rob95, How98]. educators [Bas97]. Edward [Ano95d]. Edwin [Ano91g, Hir98a]. Efficient [YH97]. Eichler [Kno90]. eigenvalue [Sar95]. Eighth [Bas98]. Eighth-Grade [Bas98]. Eilenberg [BCF^{+}98]. Electronic [Ste96]. Ellis [BC92]. emerging [Hen91]. employment [Sad97]. Encouraging [Jac90]. end [ADR97]. Ennio [Emm97, LM97]. enrollment [Run97]. entropy [LY98]. Environments [Grä94a]. enzymes [Sum95]. episode [Sau99]. equations [Way97]. era [Jac98a]. Erdos [BPV98, BS98]. Erratum [Kos95a]. Escher [Har91]. Essay [Var99]. Ethnomathematics [D’A92]. Eugene [KS98, Wig95]. Euler [Kna96a]. Evans [Ano91l]. examples [Wei96b]. experience [Ano92e]. experiences [CVG94]. Experimentation [EL95]. experiments [Pic91]. exploratorium [Pal99]. exponentiation [Mar96]. external [Wei96b]. Ezra [BBG^{+}99].

F. [Ano90i, Whi97]. Fabes [KS98]. facing [Bro97]. factorizations [ABT97]. faculty [Max97, Rob95, Run97]. Faddeev [Ano90g]. fall [Max97, Run97]. falling [GY98]. family [Mac97b]. Faris [MSV96]. Features [Ano90k]. Fefferman [Ano92c]. Fellowship [Ano94b]. Fermat [Fal95, Gra97, Jac94c, Mau97, Rib93]. few [Con95]. Fields [LLMM99, Los97, Jac92, Jac98b, LED^{+}94]. films [Ber98]. Finally [Grä94b]. Fine [And95]. finite [Ade97, Lam98a, Lam98b, SX95, Sol95]. Finsler [Che96]. first [Bro97, DMR98]. fiscal [Ano98b]. Fischer [Chi96]. five [Bor98]. Flaschka [Ano95b]. flatness [Tor97]. Floer [HWZ91]. Floyd [Ano91g]. fluid [FY99]. fluids [Con95]. focuses [Jac90]. Foias [Ano95b]. follows [Rib93]. Fonts [Grä93b]. football [Boa97]. Footer [Grä94c]. former [Ano93f, Ano93i]. forms [Duk97]. forty [Jac99c]. found [Mac98]. foundation [Ril98]. foundations [Ray99]. Four [Mil90, Run97, Tho98]. four-color [Tho98]. four-year [Run97]. Fourier [Pin95, Wal97]. Fractal [Bar96]. fractals [Str99]. FrameMaker [Dar93]. Francisco [Ano91a, Ano91b]. Frank [Ano95a]. Fred [Mac97b]. Free [Cas93]. Freudenthal [Vel91]. friend [Hir98b]. friendship [Car99]. Fuchs [ADS98]. full [Dar99]. Fulton [Ano96b]. future [Ano97e, Bab97, Bil91]. FY [Tho97].

G [Bae94, Hor94]. Galois [Kos95b, Mag99, Kos95a]. gap [Jac97a]. Garrett [Lan97b, Mac97c, You97a]. Gauge [Kot95]. Gel’fand [Ano94b]. generalization [Mau97]. genes [Lan95a, Lan95b]. genomes [Lan95a, Lan95b]. Geometric [Gar95]. geometry [Che96, Lan95c, Les97, McD98, Ray99, Tei95, Zwo99, Ano91a, Tor97, JM97, Rad96]. Georges [Bot91]. Getting [Hop95, VF97]. Giorgi [Emm97, LM97]. Givens [Poo93, Poo94]. good [Roi91]. Gorenstein [Asc92]. Goro [Ano96b]. got [Par96]. Göttingen [Seg96, Mac95a]. Gowers [Jac98b]. GP [Gou91]. Grace [Ano92f]. Grade [Bas98]. Graduate [Ano92g]. graph [Kos95a, Kos95b]. Graphics [Mar94]. Grigorievich [Goh90]. Gripenberg [CVG94]. Group [Kna96a, Kna96b, Ser97]. Groupoids [Wei96b]. Groups [Col97, Ade97, Lam98a, Lam98b, Sol95]. guide [LY98]. Gustaf [CVG94].

H [Ber90]. H. [Duf97]. Hall [Zas90].

Hamming [Mor98]. Han [DPR^{+}98]. Hans [RA96, Vel91]. Hans-Joachim [RA96]. Happen [AMS90]. Happening} [Jac94a]. Harbater [Ano95a]. hardship [Kob91]. harmonic [Kna96a, Kna96b]. Hay [Bal90]. Header [Grä94c]. Heinrich [ADS98]. Henri [Jac99e]. Henry [Bri93, Hir98a]. her [Ano91h, BHL^{+}91]. heredity [Lan95a, Lan95b]. Herrmann [Ano95b]. Hersh [Dub99]. Herz [CFV96]. Hestenes [Ano91f, Ano91i]. hiatus [Run97]. hidden [Sum95]. high [Ano97d]. him [Shi99]. history [Blu91, D’A92, Lan95d, Sau99, TW99, Kap93]. HIV [Kir96]. Hopper [Ano92f]. Horgan [Kra94]. horizon [Gra97]. Hubbard [Jer99]. hundred [Lam98a, Lam98b]. hypersurfaces [Jac95].

ICM [GJM^{+}90, Jac94b, LED^{+}94]. ICM-90 [GJM^{+}90]. ICM-94 [Jac94b, LED^{+}94]. icosahedron [Kos95a, Kos95b]. IHÉS [Jac99c]. II [Grä94a, Jac97c, Kna96b, Lam98b, Lan95b]. III [Grä94b]. image [Bar96]. immigration [Sad97]. immortality [Bae94, Hor94, Kra94]. immune [Kir96]. Implementations [Grä93b, Grä94a, Grä94b]. implications [D’A92]. IMS [Dav97b, DMR98]. incompressible [Con95]. increases [Run97]. Independent [Sos97]. infinite [Boa97]. infinity [SX95, YH97]. Instabilities [FY99]. installment [CVG94]. Institute [Gho97, Jac99a, Jac92, Jac98a, Jac99g, Jac99b]. Institution [Jac99f]. institutions [Mac93]. integrals [Ste98].

Integrated [Grä94a]. intellectual [Oke96]. interacting [Los97]. interface [Hag96]. interfaces [Fin99]. internal [Wei96b]. international [Par96]. Internet [WP98]. interpretation [Far96a]. Interview [DS97, Emm97, JK98, Jac99d, Jac99e, Jac97d, Lui97]. Interviewing [HJT98]. introduction [Ser97]. inversion [Pin95]. invests [Jac98e]. Irving [BBG^{+}99]. issue [Ano93i]. issues [Oke96, Sad97]. IV [Grä94c]. Iwasawa [Coa99].

J [Ano90d, Coa96, Whi97]. J. [Kra94]. Jackson [RA96]. Jacob [Cha98]. Jaffe [Jac97d, Jac99d]. James [CP99, Poo93, Poo94]. Jean [Ano95d]. Jean-Pierre [Ano95d]. Jerison [GH96]. Jessie [Ano90i]. Joachim [RA96]. job [HJT98]. jobs [VF97]. Johannes [ADS98]. John [Ano93e, Ano95d, CVG94, Mil98]. journals [Bab97, Ste96]. Jr [Fis92, Poo94, Zas90]. Jr. [Poo93, Whi97]. Julia [Rei96]. just [Che96].

Kac [Dol95]. Karl [Kas96]. Karp [SB96]. Keith [CHM94]. Kemeny [Ano93e]. Kenkichi [Coa99]. Kerosinka [Sau99]. keys [YH97]. Kiiti [AGHZ97]. Kleene [Kei94]. knew [Shi99]. Knuth [Ano96c]. Kodaira [Hir98b, Spe98]. Kolchin [BC92]. Konstantinovich [Ano90g]. Kontsevich [Jac98b]. Krantz [Bae94, Hor94]. Kreın [Goh90]. Kruskal [AGS94]. Kuga [Law90]. Kunihiko [Hir98b, Spe98]. Kyoto [Ano94e, Ano96c].

Lajos [DDH^{+}98]. Lamberto [Ano90e, Ano90j]. Lane [Seg96]. Langlands [Kna96a]. language [You97b]. Lars [BEH^{+}98, GKO98]. Last [Gra97, Jac94c, Mau97, Rib93, Fal95, Kos95a, Kos95b]. later [AV89, AV90]. lattice [Rot97a]. Laurent [Cha98, Cha98]. Law [LY98]. layer [Ni98]. learning [Art99]. Leendert [DS97]. Leffler [Jac99b]. Lehmer [Bri93]. Leon [GHS97, Ano94c]. lessons [Rot97b]. Letter [Knu98, Kos95a, Kos95b, Lan95c]. level [Art99]. Liang [WCA^{+}96]. librarian [BC98]. libraries [ADR97]. Lichnerowicz [BBCB^{+}99]. Life [Mac98, Jac97a, Mac97b]. Lifting [Sum95]. Lighthill [CP99]. limits [Uhl95]. Linear [Mir99]. lines [Ber98, Goo98]. Links [Eck99]. Lipman [MAC^{+}95b]. liquid [Ber98]. live [Kot95]. lives [Ano91h, BHL^{+}91, Hen91, Rot97a]. localization [Far95]. logician [Imm95]. long [Car99, Kot95]. look [CVG94, Gou91]. Lorch [Ano90h]. Louis [Ano96d, CKK^{+}98]. Louise [Bal90]. Lover [Mac97b]. Lyttle [Ano91b, Ano93b, Ano95c].

M [Duf97]. M. [Dev92a]. MA [Jer99]. MAA [DMR98, Dav97b]. Mac [Seg96]. MacArthur [Ano91j, Ano94b]. Macintosh [SHS93]. MacPherson [Ano93g]. Macsyma [CVG94]. MacWilliams [Ano90i]. Magic [Wit98]. Magnus [Ano90n, Ano91f, Ano91i]. makes [Jac90]. Manin [Ano94d]. Manuscript [AMS90]. many [Rot97a]. map [Lan95a, Lan95b]. Maple [CVG94]. Mapping [Lan95a, Lan95b]. Mario [CVG94]. Marjorie [Rad96]. Mark [Goh90]. market [Dav97a]. markup [You97b]. Marshall [Zas90]. Martin [AGS94]. Mary [MT99]. material [YH97]. Math [Bae94, Hor94, Jac90, Seg96, Sos93, Jac97b, Jac97c]. Mathematica [Dev92a, DW92]. Mathematical [Ano91j, Bab97, Gho97, Jac99f, Man95, Mar94, Saa95, Tho97, Zei93, Ano92e, Bil91, DMR99, Hop95, LLMM99, Mac93, Pal99, You97b, Jac97a, Woo90]. mathematician [Jac99b, Mac98, Hir98b, Var99]. Mathematicians [Bas97, Bas98, Dav97a, Ano91h, BHL^{+}91, Jac98e, Sad97]. mathématicien [Cha98]. Mathematics [ADR97, Ano91m, Ano93i, CHM94, CVG94, Dev92b, Dev93, Dev94a, Dev94b, Dev94c, Dev94d, Dev94e, Ewi96, GY90, Hel97, Jac97d, Jac99g, Kob91, Lew91, LS97, Mac95a, Mor99, Pic91, Sch91, Ano90l, Ano93f, Ano93g, Ano97d, Ano97e, Arn93, Art99, BPV98, Bar90, Ber98, Bro97, Chi96, Coa96, D’A92, Dau95, Dub99, EL95, GKO98, Gro98, Hag96, Hen91, Jac91, Jac97b, Jac97c, KK93, Kir96, Lan95c, Mac97b, Pal99, Par96, Pri96, Ril98, Rob95, Roi91, Run97, Sau99, Ste96, Whi97, WP98, You97a, Ano90a, Ano91b, Ano95b, Ano95c, Blu91, How98, Jac99a, Max97, TW99, Seg96]. MathType [Har93]. MATLAB [DDW93]. matrix [BP99, Van97, Wit98]. matter [Los97]. Maurice [PR95]. MAX [CVG94, Ewi93, Jac98a]. McMullen [Jac98b]. measures [Tor97]. Measuring [CW98]. mechanics [Van97, Far96a]. Medal [AGS94, Ano92b, Ano96d]. Medalists [LLMM99]. Medals [Ano91k, SB96, Jac98b, LED^{+}94]. meets [WP98]. Memorial [Ano94c]. memoriam [LM96]. Memories [MSV96, Car99]. Menger [Kas96]. Merging [Hen91]. methods [YH97]. Meyer [GH96]. MG [Mar94]. Michael [CP99]. Michel [Ano95a]. Michio [ABFS99, Law90]. Microsoft [Jac98e]. millions [KR97, Mum97]. Mina [GLMM98]. mind [Far96b, Mil98]. Misdemeanours [Sto91]. Missing [Eck99, Far96b]. Mittag [Jac99b]. Mittag-Leffler [Jac99b]. Model [Mar96]. models [Lan95a, Lan95b]. modern [Lan97b, Mac97c, Woo90, Mac97d]. Moduli [Uhl95]. Moishezon [Ano93d]. Montgomery [Bor92]. month [CVG94]. Moody [Dol95]. Mordell [Lan95c]. Mori [Ano90m]. Morita [AGHZ97]. Morris [Ber90]. Moscow [Sos97]. motion [FY99]. moving [Ber98]. MR1139519 [Var99]. MR1254647 [Kra94]. MR1291602 [Far96a]. MR1303779 [JM97]. MR1340198 [Rad96]. MR1350011 [Seg96]. MR1462890 [Dub99]. MR1625902 [Jer99].

MR1631630 [Mil98]. Much [Jac97a]. Murray [Ano92f]. mystery [Wit98].

NAS [Ano93g, Coa96]. Nasar [Mil98]. Nash [Mil98]. Nathan [And95]. national [Par96, AGS94, Ano91k, Ano92b, Ano96d, Bas98, SB96]. Nazis [Mac95a, Seg96]. NCTM [How98]. need [OP97]. needed [Jac97a]. needs [Roi91]. Nelson [Ano95a, Ano95d]. Nemmers [Ano94d]. Neugebauer [Boa90]. Nevanlinna [Jac98d, LED^{+}94]. Newsom [Pri90]. Newton [AV89, AV90]. Nicolas [Bor98]. Niels [Var98]. Nirenberg [Ano96d]. no [Bae94, Hor94, Kra94, Seg96]. Noncommutative [Les97, JM97]. nonlinear [Way97]. nonpositive [Kno90]. Norbert [Ano90a, Ano95b, JS95, Man95]. Notices [Bae94, Hor94, Seg96]. NSF [Ano98b]. Nucci [Dev92a].

Observations [BB92]. Odile [Cha98]. Off [SX95]. Offered [Ano98a]. old [Duk97, Jac97a]. old-time [Jac97a]. Olga [LM96]. Omnès [Far96a]. open [Con95]. opens [Jac92]. opinions [Col97]. opportunities [Ano90l]. Optional [Pri96]. Orono [Ano91c]. Ostrowski [Ano90f]. Oswald [Ano91a]. Otto [Boa90]. overview [Par96]. own [Ano91h, BHL^{+}91]. Oxford [Dub99].

P [Ano92b, Fis92]. Pacific [Gho97].

package [Cas91]. Paper [AMS90]. PARI} [Gou91]. Paris [Cha98]. Part [AMS90, CVG94]. partial [Way97]. partnership [Rob95]. past [Bab97, Bil91]. Paul [Wig95, BPV98, BS98, Bar97]. pedagogical [D’A92]. Pedersen [Var98]. Pedro [CKSW98]. Periodic [Way97]. Perlis [Ano90d]. Perrin [Ano99b]. Perrin-Riou [Ano99b]. personal [Bab97]. perspective [BC98]. perspectives [Blu91, Sch91]. Peter [Jac98d]. Peters [Jer99]. philosophy [D’A92]. physics [Col97, Dol95, Zwo99]. Pick [Oss99]. Pierre [Ano95d]. pipeline [Lew91]. plan [Ano90l]. Planck [Jac98a]. plane [Mir99]. player [Boa97]. Please [Tur91]. Pocket [Leh94]. Poincaré [Kno90]. point [Bes98]. Pointwise [Pin95]. Poland [Hel97]. popular [Sos93]. Possible [Gro98]. postdoctoral [Ano92e]. potential [Ano97e]. power [Woo90]. precedes [Dev92a]. prerequisites [OP97]. present [Bab97, Bil91]. presented [Ano91d, LED^{+}94]. presents [CVG94]. President [Jac97d, Jac99d]. Presidential [Mor99]. Presidents [Blu91]. Press [Dub99, JM97, Rad96]. Price [Zei93]. Pricing [Bes98]. prime [ABT97]. Principia [AV89, AV90]. prises [Cha98]. Prize [Ano95a, Ano95b, Ano95c, Ano94d, Ano99a, Ano99b, Mau97, Ano90f, Ano90a, Ano90m, Ano91a, Ano91b, Ano92c, Ano92a, Ano92d, Ano93b, Ano93h, Ano94c, Ano94e, Ano96c, Ano97a, Ano97b, Bar97, Jac98d, LED^{+}94, ST94]. Prizes [Ano90b, Ano90c, Ano91e, Ano91c, Ano93a, Ano93c, Ano94a, Ano95d, Ano96a, Ano96b, Ano97c, Ano99c]. probabilistic [Lan95a, Lan95b]. probe [Sum95]. Problem [Lan97a, Mau97]. problems [Con95, Duk97, Jac90]. Processors [Mil90, Mil92]. producers [Jac91]. profit [Jac98e]. program [D’A92, Gra97, Jac98c]. progress [Sch91]. prologue [Kna99]. proof [Dar99, EL95, Fal95, Jac94c, Kra94, Bae94, Hor94, Kra94]. property [Oke96]. proves [Rib93]. proving [Ano91d]. Pseudorandomness [Gol99]. Publication [Jac94a]. Publications [Ano98a, Bes98]. Published [AMS90]. publisher [Bes98]. publishing [BC98]. Pukánszky [DDH^{+}98].

Quadratic [Kap93, Che96, Duk97]. quantum [Far96a, Van97]. quarters [Sos97]. Quartic [Leh94]. Quasicrystals [Rad96]. Questionnaire [Ano93f]. questions [Art99].

R [Ano90h, Ano91f, Ano96b]. R.

[Dub99, Fal95]. Rademacher [Kno90]. Ralph [Fis92]. Random [ABT97, Far95]. Raynaud [Ano95a]. read [AV89, AV90]. Real [Jac95, Tur91]. realities [LS97]. really [Dub99]. receive [Ano91j, Ano92d, Jac98b, SB96, ST94]. Receives [Ano96c, AGS94, Ano90f, Ano92b, Ano93g, Ano94b, Ano94c, Ano94d, Ano94e, Ano96d, Coa96, Jac98d]. reciprocity [Kap93]. Rees [GLMM98]. reform [Jac97b, Jac97c, KR97, Mum97]. Reforming [BC98]. regarding [Con95]. related [Ano90l]. Relations [Eck99]. Remembering [MAC^{+}95b]. reminiscences [CDK^{+}95]. Renewing [Ano90l]. report [Dav97b, DMR98]. reports [CVG94]. representation [Rei97]. Representations [Lam98a, Lam98b, Kna96a, Kna96b]. request [Ano98b]. Requiem [Mac97a]. rescue [Mor99]. research [ADR97, Ano90l, Art99, D’A92]. researchers [Oke96]. Resonances [Zwo99]. responds [Hor94]. Response [Bae94]. restriction [Che96]. results [Con95, Duk97]. retiring [Mor99]. Review [Dub99, Far96a, Gra97, Jer99, JM97, Cha98, CVG94, DDW93, Lan95c, Rad96, Var99]. reviews [CVG94, Dev92a, Jac97a, Woo90]. revision [How98]. Rham [Bot91, Eas99]. Richard [Mor98]. Riemannian [Che96]. Rigor [Gil97]. Rigorous [Zei93]. Riou [Ano99b]. Robert [Ano93g, Wei96a]. Robinson [Rei96]. Roland [Far96a, MSV96]. roles [Ste98]. RSA [Bon99]. Russian [Arn93, Mac93, MSV96, Sos93]. Ruth [Ano91b, Ano93b, Ano95c].

S [Ano96b, Duf97]. S.

[Ano96b, Bae94, Hor94, Mil98, Seg96]. Sah [DPR^{+}98]. Samuel [BCF^{+}98]. San [JM97, Ano91a, Ano91b]. Sarvadaman [AHW98]. Satter [Ano91b, Ano93b, Ano95c, Ano97b, Ano99b]. scholarly [BC98]. Scholars [Jac98c]. school [Ano97d, Lew91, Rob95]. Schuster [Mil98]. Schwartz [Cha98, Cha98]. Schwarz [Oss99]. Sci. [Kra94]. Science [Jac90, Far96b, AGS94, Ano91k, Ano92b, Ano96d, SB96]. sciences [Bil91, BC98, DMR99, Tho97, Gho97]. Scientific [Mil92, Bes98]. scientists [Ano91j, Ano92e]. search [Far96b]. second [Ano93g, CVG94, Dav97b, LY98]. Segal [BBG^{+}99]. Selberg [Sar95]. Semi [Zei93]. Semi-Rigorous [Zei93]. Senechal [Rad96]. senior [Col97]. series [Boa97, Kno90]. Serre [Ano95d]. set [Goo98]. several [Pin95, Pic91]. Shadows [Far96b]. Shanks [Wil97]. shape [CW98]. Shen [JK98]. Shigefumi [Ano90m]. Shiing [JK98]. Shimura [Ano96b, Dar99, Lan95d]. Shor [Jac98d]. short [Pic91]. siècle [Cha98]. Siegel [Lan95c]. sieves [Pom96]. sign [BP99]. Simon [Ano94c, Mil98]. simple [Saa95, Sol95]. Single [Doo90]. Singular [Ste98]. sister [Rei96]. Siu [Ano93h]. six [Ano91h, BHL^{+}91]. sketch [RA96]. skillful [Mac97a]. sky [GY98]. Smale [SB96]. Smithsonian [Jac99f]. Soc [Bae94, Hor94, Seg96]. societies [Lan97a]. software [CVG94]. solutions [Way97]. solved [BP99]. solving [Lan97a]. Some [BB92, Duk97, Lan95d, Mac93, CDK^{+}95, Wei96b]. Southern [LS97]. Soviet [Ano93f, Ano93i, KK93, Sau99]. space [Goo98]. spaces [Tei95, Uhl95]. Spanier [Hir98a]. Special [Ano93i]. Spiegel [GLMM98]. spike [Ni98]. spike-layer [Ni98]. Spreadsheets [CVG94]. stability [Los97]. staircase [Har91]. Stand [Tur91]. standards [How98]. started [Hop95]. state [Ril98]. states [Ni98, Ano92e]. steady [Ni98]. Steele [Ano90b, Ano91c, Ano93a, Ano93c, Ano94a, Ano95d, Ano96a, Ano96b, Ano97c, Ano99c]. Stephen [Kei94]. still [Roi91]. strong [Ril98]. Stroock [Ano96b]. structures [ABT97, McD98]. Student [DDW93]. study [Ano92e]. Sun [Ano95c]. Sun-Yung [Ano95c]. surface [Fin99]. survey [Dav97b, DMR99, Lan97b, Run97, DMR98, Mac97c]. survive [Arn93]. Suzuki [ABFS99]. Swedish [Jac99b]. Symmetric [FG95, Tei95]. Symmetry [Rad95, Dol95, Wei96b]. Symplectic [McD98]. System [Mar94, Dev92a]. systems [Kup99, Mir99].

T [Ano95d]. Takes [Jac94a, CVG94]. tale [Pom96]. Taniyama [Dar99, Lan95d, Rib93]. Tate [Ano95d]. taught [Rot97b]. Taussky [LM96]. Taussky-Todd [LM96]. Taylor [Fal95]. Teach [Chu92, Knu98, RS99]. teacher [Hir98b]. teaching [Art99]. technology [Woo90]. Ten [Rot97b]. Test [Bas98]. their [Ano91h, BHL^{+}91, Ni98, Sol95]. theorem [Ano91d, Fal95, Tho98, Gra97, Jac94c, Mau97, Rib93]. Theorems [Zei93]. Theory [Jac98e, Kot95, Mag99, Mar96, Rei97, Rot97a, Ser97]. Thermodynamics [LY98]. thin [Ber98]. Third [Kob91]. Thomason [Wei96a]. Thompson [Ano92d]. Three [CVG94, Mil92]. tilings [Rad95]. time [Jac97a, SX95]. Today [Ano97e]. Todd [LM96]. together [Rob95]. tomography [Gar95]. Tomorrow [Zei93, Cas91]. Top [Jac91]. Topology [Eck99, Bor99, Sum95]. tour [Wei96b]. Trade [Sto91]. traditions [Sos93]. transition [Ano92g]. Translated [MSV96]. treasures [Jac99f]. trends [Gro98]. Trevor [Ano91l]. tribute [SW96]. truncated [Kos95a, Kos95b]. Tucker [CDK^{+}95]. twentieth [ADR97]. Twenty [Bon99, Bor98, Bro97]. twenty-first [Bro97]. Twenty-five [Bor98]. two [CVG94, Dev92a, Pom96]. two-part [CVG94]. typesetting [Cas91].

U.S. [Ano90l]. uncertain [Ano97e]. uncertainty [Chi96]. undergraduate [Lew91]. understand [Kir96]. unified [Van97]. unifying [Wei96b]. Union [Ano93f, Ano93i]. United [Ano92e]. Univ [Dub99, Rad96]. universe [CW98]. University [Sos97, Art99, Rob95, Run97]. Update [Jac94c, Tho98]. Using [Chu92, Kir96, Lan95a, Lan95b, Sum95].

Valentine [LSW90]. Valerian [BEH^{+}98, GKO98]. Varadhan [Ano96b]. variables [Pin95]. Variations [Eas99]. Vassallo [CVG94]. Veblen [Ano91a]. Version [Grä94b, Dev92a]. view [Bab97, Bes98]. Views [Ano97d]. Vigand [Var98]. Vincent [Pri90]. visualization [Pal99]. Vladimir [Lui97]. VTEX [Cas91].

W [CDK^{+}95, Wei96a]. Waerden [DS97, Mac97d]. Wallace [Poo93, Poo94]. war [Hel97]. Warren [SW96]. Wars [Jac97b, Jac97c]. was [BP99]. wave [Van97]. wavelet [Wal97]. Wavelets [Dau95, Jer99]. waves [Far95]. wealth [Ano97e]. Wei [WCA^{+}96]. Wei-Liang [WCA^{+}96]. weights [Kno90]. Weil [Ano94e, BCC^{+}99, Bor99, Car99, Dar99, Kna99, Ray99, Shi99, Var99]. Wellesley [Jer99]. Wesley [Mor98]. Where [WP98, Par96]. Whiteman [GHS97]. Whose [Oke96]. Wiener [Ano90a, Ano95b, JS95, Man95]. Wigner [Wig95]. Wiles [Coa96, Fal95, Rib93]. Wilhelm [Ano90n]. Will [Arn93, Tur91]. William [Ano96b, MSV96]. wins [Ano93h]. wish [Rot97b]. without [Che96]. Wolf [Ano90c, Ano92d]. Wolfgang [ADS98]. Wolfskehl [Bar97]. Women [Ano91m, Blu91, Jac90, KK93, TW99, Bil91, Dau95, Jac91, Lew91, Roi91, Sch91, VF97, Hen91]. Word [Mil90, Mil92]. words [Ano91h, BHL^{+}91]. work [LLMM99, Man95]. working [Rob95]. Workshops [Chu92]. world [Jer99, Kob91]. Writing [Hwa95].

Yau [ST94]. year [Ano98b, Run97]. Years [Bon99, AV89, AV90, Bor98, Lam98a, Lam98b]. York [Dub99, Mil98]. Young [Jac98c]. Yung [Ano95c].

Zorn [Ewi93]. Zürich [Jac94b, LED^{+}94]. Zygmund [CS92, Ste98].

References

Aschbacher:1999:MS

[ABFS99] Michael Aschbacher, Helmut Bender, Walter Feit, and Ronald Solomon. Michio Suzuki (1926– 1998). Notices of the American Mathematical Society, 46(5):543– 551, May 1999. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arratia:1997:RCS

[ABT97] Richard Arratia, A. D. Barbour, and Simon Tavaré. Random combinatorial structures and prime factorizations. Notices of the American Mathematical Society, 44(8):903–910, September 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                                   tronic). URL https://www.ams.
                                                   org/notices/199708/tavare.
                                                   pdf.

Adem:1997:RDC

[Ade97] Alejandro Adem. Recent developments in the cohomology of finite groups. Notices of the American Mathematical Society, 44(7):

806–812, August 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
notices/199707/adem.pdf.

Anderson:1997:MRL

[ADR97] N. D. Anderson, K. Dilcher, and J. Rovnyak. Mathematics research libraries at the end of the twentieth century. Notices of the American Mathematical Society, 44(11):1469–1472, December 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199711/
comm-rovnyak.pdf.

Anderson:1998:WHJ

[ADS98] J. Milne Anderson, David Drasin, and Linda R. Sons. Wolfgang Heinrich Johannes Fuchs (1915– 1997). Notices of the American Mathematical Society, 45 (11):1472–1478, December 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.
ams.org/notices/199811/mem-
fuchs.pdf.

Arhangelskii:1997:KM

[AGHZ97] A. V. Arhangel’skii, Kenneth R.

Goodearl, and B. Huisgen-Zimmermann. Kiiti Morita (1915–1995). Notices of the American Mathematical Society, 44(6):680–684, June/July 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.

                                        org/notices/199706/morita.
                                        pdf.

Ablowitz:1994:MDK

[AGS94] Mark J. Ablowitz, John Greene, and Harvey Segur. Martin D. Kruskal receives National Medal of Science. Notices of the American Mathematical Society, 41 (3):182–184, March 1994. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ayoub:1998:SC

[AHW98] Raymond G. Ayoub, James G. Huard, and Kenneth S. Williams. Sarvadaman Chowla (1907– 1995). Notices of the American Mathematical Society, 45(5): 594–598, May 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL https://www.ams.org/
                                        notices/199805/comm-chowla.
                                        pdf.

AMSCSD:1990:MPP

[AMS90] AMS Composition Services Department. From manuscript to published paper: How does it happen? part 2. Notices of the American Mathematical Society, 37(8):1027–1028, October 1990. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Andrews:1995:NF

[And95] George E. Andrews. Nathan Fine (1916–1994). Notices of the American Mathematical Society, 42(6):678–679, June 1995.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199506/
fine.pdf.

Anonymous:1990:NWP

[Ano90a] Anonymous. 1990 Norbert Wiener Prize in Applied Mathematics awarded in Columbus. Notices of the American Mathematical Society, 37(7):808–811, September 1990. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:SPA

[Ano90b] Anonymous. 1990 Steele Prizes awarded in Columbus. Notices of the American Mathematical Society, 37(7):801–807, September 1990. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:WP

[Ano90c] Anonymous. 1990 Wolf Prizes.

Notices of the American Mathematical Society, 37(3):286, March 1990. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:AJP

[Ano90d] Anonymous. Alan J. Perlis 1922– 1990. Notices of the American Mathematical Society, 37(4): 441, April 1990. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:CL

[Ano90e] Anonymous. Cesari Lamberto 1910–1990. Notices of the American Mathematical Society, 37(6): 677, July/August 1990. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:BRO

[Ano90f] Anonymous. de Branges receives Ostrowski Prize. Notices of the American Mathematical Society, 37(6):678–679, July/ August 1990. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:DKF

[Ano90g] Anonymous. Dmitriĭ Konstantinovich Faddeev: 1906–1989. Notices of the American Mathematical Society, 37(6):677–678, July/ August 1990. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:ERL

[Ano90h] Anonymous. Edgar R. Lorch: 1907–1990. Notices of the American Mathematical Society, 37(6): 678, July/August 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:FJM

[Ano90i] Anonymous. F. Jessie MacWilliams: 1917–1990. Notices of the American Mathematical Society, 37 (8):1033, September 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:LC

[Ano90j] Anonymous. Lamberto Cesari:

1910–1990. Notices of the American Mathematical Society, 37(6): 677, July/August 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:NTF

[Ano90k] Anonymous. New TEX features from AMS. Notices of the American Mathematical Society, 37(2): 143–144, February 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:RUM

[Ano90l] Anonymous. Renewing U.S.

mathematics: a plan for the 1990s. Appendix B: Recent research accomplishments and related opportunities. Notices of the American Mathematical Society, 37(8):984–1008, October 1990. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With a commentary by Allyn Jackson.

Anonymous:1990:SMA

[Ano90m] Anonymous. Shigefumi Mori awarded 1990 Cole Prize in Algebra. Notices of the American Mathematical Society, 37(2): 118–119, February 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1990:WM

[Ano90n] Anonymous. Wilhelm Magnus: 1907–1990. Notices of the American Mathematical Society, 37 (10):1372, November 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1991:OVP

[Ano91a] Anonymous. 1991 Oswald Veblen Prize in Geometry awarded in San Francisco. Notices of the American Mathematical Society, 38(3):181–183, March 1991. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1991:RLS

[Ano91b] Anonymous. 1991 Ruth Lyttle Satter Prize in Mathematics awarded in San Francisco. Notices of the American Mathematical Society, 38(3):185–187, March 1991. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1991:SPA

[Ano91c] Anonymous. 1991 Steele Prizes awarded in Orono. Notices of the American Mathematical Society, 38(8):884–890, September 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1991:ATP

[Ano91d] Anonymous. Automatic theorem proving awards presented.

Notices of the American Mathematical Society, 38(5):405–410, May/June 1991. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1991:BPA

[Ano91e] Anonymous. Bergman Prizes awarded. Notices of the American Mathematical Society, 38(7): 792–793, August 1991. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1991:DMR

[Ano91f] Anonymous. Deaths: Magnus R.

Hestenes. Notices of the American Mathematical Society, 38(6): 657, July/August 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
journals/notices/199107.

Anonymous:1991:EF

[Ano91g] Anonymous. Edwin Floyd: 1924– 1990. Notices of the American Mathematical Society, 38(2):113, February 1991. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1991:HOW

[Ano91h] Anonymous. In her own words:

six mathematicians comment on their lives and careers. Notices of the American Mathematical Society, 38(7):702–706, August 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Women in mathematics.

Anonymous:1991:MH

[Ano91i] Anonymous. Magnus Hestenes: 1906–1991. Notices of the American Mathematical Society, 38(7): 792, September 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL https://www.ams.org/
                                        journals/notices/199109.

Anonymous:1991:MSR

[Ano91j] Anonymous. Mathematical scientists receive MacArthur Awards. Notices of the American Mathematical Society, 38(7): 793, August 1991. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1991:NMS

[Ano91k] Anonymous. National Medals of Science. Notices of the American Mathematical Society, 38(4):277– 284, April 1991. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1991:TE

[Ano91l] Anonymous. Trevor Evans: 1925–1991. Notices of the American Mathematical Society, 38(7): 792, August 1991. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1991:WM

[Ano91m] Anonymous. Women in Mathematics. American Mathematical Society, Providence, RI, USA, 1991. ISSN 0002-9920 (print), 1088-9477 (electronic). 701–775 pp. Notices Amer. Math. Soc. 38 (1991), no. 7.

Anonymous:1992:CPA

[Ano92a] Anonymous. 1992 Cole Prize in algebra. Notices of the American Mathematical Society, 39(4):275– 277, April 1992. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1992:APC

[Ano92b] Anonymous. Alberto P. Calderón receives National Medal of Science. Notices of the American Mathematical Society, 39(4):283– 285, April 1992. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1992:BPA

[Ano92c] Anonymous. Bergman Prize awarded to Charles Fefferman. Notices of the American Mathematical Society, 39(4):320–321, April 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1992:CTR

[Ano92d] Anonymous. Carleson and Thompson receive Wolf Prize. Notices of the American Mathematical Society, 39(4):321–322, April 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1992:EMS

[Ano92e] Anonymous. Educating mathematical scientists: doctoral study and the postdoctoral experience in the United States. Notices of the American Mathematical Society, 39(5):390–397, May 1992.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With discussion.

Anonymous:1992:GMH

[Ano92f] Anonymous. Grace Murray Hopper 1906–1992. Notices of the American Mathematical Society, 39(4):320, April 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1992:GET

[Ano92g] Anonymous. Graduate education in transition. Notices of the American Mathematical Society, 39(5):398–403, May 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1993:SPa

[Ano93a] Anonymous. 1992 Steele Prizes. Notices of the American Mathematical Society, 40(3):221–225, March 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1993:RLS

[Ano93b] Anonymous. 1993 Ruth Lyttle Satter Prize. Notices of the American Mathematical Society, 40(3):229–230, March 1993. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1993:SPb

[Ano93c] Anonymous. 1993 Steele Prizes. Notices of the American Mathematical Society, 40(8):973–977, September 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1993:BM

[Ano93d] Anonymous. Boris Moishezon (1937–1993). Notices of the American Mathematical Society, 40(9):1220, October 1993. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1993:JK

[Ano93e] Anonymous. John Kemeny 1926– 1992. Notices of the American Mathematical Society, 40(3): 243, March 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1993:QMF

[Ano93f] Anonymous. Questionnaire on mathematics in the former Soviet Union. Notices of the American Mathematical Society, 40(2): 132–137, February 1993. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Special issue on mathematics in the former Soviet Union.

Anonymous:1993:RDM

[Ano93g] Anonymous. Robert D. MacPherson receives second NAS Award in mathematics. Notices of the American Mathematical Society, 40(4):325–326, April 1993. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1993:SWB

[Ano93h] Anonymous. Siu wins Bergman Prize. Notices of the American Mathematical Society, 40(5):473– 474, May 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1993:SIM

[Ano93i] Anonymous. Special issue on Mathematics in the former Soviet Union. American Mathematical Society, Providence, RI, USA, 1993. ISSN 0002-9920 (print), 1088-9477 (electronic). 103–139 pp. Notices Amer. Math. Soc. 40 (1993), no. 2.

Anonymous:1994:SP

[Ano94a] Anonymous. 1994 Steele Prizes. Notices of the American Mathematical Society, 41(8):905–912, September 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1994:GRM

[Ano94b] Anonymous. Gel’fand receives MacArthur Fellowship. Notices of the American Mathematical Society, 41(7):797–798, August 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1994:LSR

[Ano94c] Anonymous. Leon Simon receives 1994 Bôcher Memorial Prize. Notices of the American Mathematical Society, 41(2):99–100, February 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1994:MRN

[Ano94d] Anonymous. Manin receives Nemmers prize. Notices of the American Mathematical Society, 41(7):796–797, August 1994. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1994:WRK

[Ano94e] Anonymous. Weil receives Kyoto Prize. Notices of the American Mathematical Society, 41(7):793– 794, August 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1995:FNC

[Ano95a] Anonymous. 1995 Frank Nelson Cole Prize in Algebra [awarded to David Harbater and Michel Raynaud]. Notices of the American Mathematical Society, 42 (4):454–456, April 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199504/199504-
toc.html.

Anonymous:1995:NWP

[Ano95b] Anonymous. 1995 Norbert Wiener Prize in Applied Mathematics (awarded to Herrmann Flaschka and Ciprian Foias). Notices of the American Mathematical Society, 42(4):457–458, April 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199504/
199504-toc.html.

Anonymous:1995:RLS

[Ano95c] Anonymous. 1995 Ruth Lyttle Satter Prize in Mathematics [awarded to Sun-Yung Alice Chang]. Notices of the American Mathematical Society, 42 (4):459–460, April 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199504/199504-
                                        toc.html.

Anonymous:1995:SP

[Ano95d] Anonymous. 1995 Steele Prizes (awarded to Edward Nelson, Jean-Pierre Serre, and John T. Tate). Notices of the American Mathematical Society, 42 (11):1288–1292, November 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199511/199511-
                                        toc.html.

Anonymous:1996:SP

[Ano96a] Anonymous. 1996 Steele Prizes. Notices of the American Mathematical Society, 43(11):1340– 1347, November 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1996:SPB

[Ano96b] Anonymous. 1996 Steele Prizes (Bruce C. Berndt, William Fulton, Daniel Stroock, S. R. S. Varadhan, and Goro Shimura). Notices of the American Mathematical Society, 43(11):1340– 1347, November 1996. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1996:KRK

[Ano96c] Anonymous. Knuth receives Kyoto Prize. Notices of the American Mathematical Society, 43 (10):1159–1160, October 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.
org/notices/199610/people.
pdf.

Anonymous:1996:LNR

[Ano96d] Anonymous. Louis Nirenberg receives National Medal of Science. Notices of the American Mathematical Society, 43 (10):1111–1116, October 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions by Luis Caffarelli and Joseph J. Kohn.

Anonymous:1997:CP

[Ano97a] Anonymous. 1997 Cole Prize.

Notices of the American Mathematical Society, 44(3):346–347, March 1997. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1997:SPb

[Ano97b] Anonymous. 1997 Satter Prize.

Notices of the American Mathematical Society, 44(3):348–349, March 1997. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1997:SPa

[Ano97c] Anonymous. 1997 Steele Prizes. Notices of the American Mathematical Society, 44(3):342–345, March 1997. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1997:VHS

[Ano97d] Anonymous. Views on high school mathematics education. Notices of the American Mathematical Society, 44(2):197–206, February 1997. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
notices/199702/comm-views.
pdf.

Anonymous:1997:WPU

[Ano97e] Anonymous. A wealth of potential but an uncertain future: Today’s mathematics departments. Notices of the American Mathematical Society, 44 (4):439–443, April 1997. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199704/
comm-future.pdf.

Anonymous:1998:NPO

[Ano98a] Anonymous. New publications offered by the AMS. Notices of the American Mathematical Society, 45(10):1379–1391, November 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/199810/
newpubs-nov98.pdf.

Anonymous:1998:NFY

[Ano98b] Anonymous. NSF fiscal year 1999 budget request. Notices of the American Mathematical Society, 45(5):616–618, May 1998. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199805/
comm-nsfbudget.pdf.

Anonymous:1999:BP

[Ano99a] Anonymous. 1999 Bôcher prize.

Notices of the American Mathematical Society, 46(4):463–466, April 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1999:SPA

[Ano99b] Anonymous. 1999 Satter prize [awarded to Bernadette Perrin-Riou]. Notices of the American Mathematical Society, 46(4):467– 468, April 1999. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1999:SP

[Ano99c] Anonymous. 1999 Steele Prizes.

Notices of the American Mathematical Society, 46(4):457–462, April 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arnold:1993:WRM

[Arn93] V. I. Arnol’d. Will Russian mathematics survive? Notices of the American Mathematical Society, 40(2):104–107, February 1993. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic). Special issue on mathematics in the former Soviet Union.

Artigue:1999:TLM

[Art99] Michèle Artigue. The teaching and learning of mathematics at the university level. Crucial questions for contemporary research in education. Notices of the American Mathematical Society, 46(11):1377–1385, December 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199911/
199911-toc.html.

Aschbacher:1992:DG

[Asc92]  Michael Aschbacher.  Daniel

Gorenstein (1923–1992). Notices of the American Mathematical Society, 39(10):1190–1191, November 1992. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arnold:1989:NPR

[AV89] V. I. Arnol’d and V. A. Vasil’ev. Newton’s Principia read 300 years later. Notices of the American Mathematical Society, 36(9): 1148–1154, November 1989. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See addendum [AV90].

Arnold:1990:ANP

[AV90] V. I. Arnol’d and V. A. Vasil’ev. Addendum to: “Newton’s Principia read 300 years later”. Notices of the American Mathematical Society, 37(2):144, February 1990. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See [AV89].

Babbitt:1997:MJP

[Bab97] Donald Babbitt. Mathematical journals: past, present and future – a personal view. Notices of the American Mathematical Society, 44(1):29–32, January 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baer:1994:RIP

[Bae94]  Robert M. Baer.  Response

to: “The immortality of proof” [Notices Amer. Math. Soc. 41 (1994), no. 1, 10–13; 1 254 635] by S. G. Krantz. Notices of the American Mathematical Society, 41(7):740–741, August 1994. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See [Kra94].

Baldwin:1990:LH

[Bal90] John T. Baldwin. Louise Hay 1935–1989. Notices of the American Mathematical Society, 37(1): 20, January 1990. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barwise:1990:CM

[Bar90] Jon Barwise. Computers and mathematics. Notices of the American Mathematical Society, 37(1):7–8, January 1, 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barnsley:1996:FIC

[Bar96] Michael F. Barnsley. Fractal image compression. Notices of the American Mathematical Society, 43(6):657–662, June 1996. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barner:1997:PWW

[Bar97] Klaus Barner. Paul Wolfskehl and the Wolfskehl Prize. Notices of the American Mathematical Society, 44(10):1294– 1303, November 1997. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199710/barner.
                                        pdf.

Bass:1997:ME

[Bas97] Hyman Bass. Mathematicians as educators. Notices of the American Mathematical Society, 44(1):18–21, January 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199701/
                                        comm-bass.pdf.

Bass:1998:MNE

[Bas98] Hyman Bass. Mathematicians and the National Eighth-Grade Test. Notices of the American Mathematical Society, 45(5): 589–593, May 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL https://www.ams.org/
                                        notices/199805/bass.pdf.

Borwein:1992:SOC

[BB92] Jonathan Borwein and Peter B. Borwein. Some observations on computer aided analysis. Notices of the American Mathematical Society, 39(8): 825–829, October 1992. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.
org/journals/notices/199210/
199210FullIssue.pdf.

Berger:1999:AL

[BBCB^{+}99] Marcel Berger, Jean-Pierre Bourguignon, Yvonne Choquet-Bruhat, Charles-Michel Marle, and André Revuz. André Lichnerowicz (1915–1998). Notices of the American Mathematical Society, 46(11):1387–1396, December 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199911/
199911-toc.html. The segments

of this article, except for the one by Bourguignon, are translated from the French.

Baez:1999:IES

[BBG^{+}99] John C. Baez, Edwin F. Beschler, Leonard Gross, Bertram Kostant, Edward Nelson, Michèle Vergne, and Arthur S. Wightman. Irving Ezra Segal (1918–1998). Notices of the American Mathematical Society, 46(6):659–668, June/ July 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bass:1992:EK

[BC92] Hyman Bass and Phyllis Cassidy. Ellis Kolchin (1916–1991). Notices of the American Mathematical Society, 39(6):599–600, June/July 1992. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Branin:1998:RSP

[BC98] Joseph J. Branin and Mary Case. Reforming scholarly publishing in the sciences: a librarian perspective. Notices of the American Mathematical Society, 45 (4):475–486, April 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199804/branin.
                                        pdf.

Borel:1999:AW

[BCC^{+}99] Armand Borel, Pierre Cartier, Komaravolu Chandrasekharan, Shiing-Shen Chern, and Shokichi Iyanaga. André Weil (1906– 1998). Notices of the American Mathematical Society, 46(4):440– 447, April 1999. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bass:1998:SE

[BCF^{+}98] Hyman Bass, Henri Cartan, Peter Freyd, Alex Heller, and Saunders MacLane. Samuel Eilenberg (1913–1998). Notices of the American Mathematical Society, 45(10):1344–1352, November 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199810/
mem-eilenberg.pdf.

Bott:1998:LVA

[BEH^{+}98] Raoul Bott, Clifford Earle, Dennis Hejhal, James Jenkis, Troels Jorgensen, Steven G. Krantz, Albert Marden, and Robert Osserman. Lars Valerian Ahlfors (1907–1996). Notices of the American Mathematical Society, 45(2):248–255, February 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199802/
comm-krantz.pdf. Edited and

with an introduction by Steven G. Krantz.

Berger:1990:MHD

[Ber90] James Berger. Morris H. DeGroot, 1931–1989. Notices of the American Mathematical Society, 37(3):285–286, March 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bertozzi:1998:MMC

[Ber98] Andrea L. Bertozzi. The mathematics of moving contact lines in thin liquid films. Notices of the American Mathematical Society, 45(6):689–697, June/July 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Beschler:1998:PSP

[Bes98] Edwin F. Beschler. Pricing of scientific publications: a commercial publisher’s point of view. Notices of the American Mathematical Society, 45 (10):1333–1343, November 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199810/
beschler.pdf.

Birman:1991:HOW

[BHL^{+}91] Joan S. Birman, Deborah Haimo, Susan Landau, Bhama Srinivasan, Vera S. Pless, and Jean E. Taylor. In her own words. Six mathematicians comment on their lives and careers. Notices of the American Mathematical Society, 38(7):702–706, August 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Billard:1991:PPF

[Bil91] Lynne Billard. The past, present, and future of academic women in the mathematical sciences. Notices of the American Mathematical Society, 38(7):707–714, August 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Women in mathematics.

Blum:1991:BHA

[Blu91] Lenore Blum. A brief history of the Association for Women in Mathematics: the Presidents’ perspectives. Notices of the American Mathematical Society, 38(7):738–754, August 1991. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Women in mathematics.

Boas:1990:ON

[Boa90] R. P. Boas. Otto Neugebauer:

1899–1990. Notices of the American Mathematical Society, 37 (5):541, May/June 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Boas:1997:FPI

[Boa97] Harold P. Boas. The football player and the infinite series. Notices of the American Mathematical Society, 44(11):1430– 1435, December 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
notices/199711/boas.pdf.

Boneh:1999:TYA

[Bon99] Dan Boneh. Twenty years of attacks on the RSA cryptosystem. Notices of the American Mathematical Society, 46(2): 203–213, February 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://theory.
stanford.edu/~dabo/papers/
RSA-survey.pdf;   http://theory.
stanford.edu/~dabo/papers/
RSA-survey.ps;   https://www.
ams.org/notices/199902/199902-
toc.html.

Borel:1992:DM

[Bor92] Armand Borel. Deane Montgomery (1909–1992). Notices of the American Mathematical Society, 39(7):684–686, August 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Borel:1998:TFY

[Bor98] Armand Borel. Twenty-five years with Nicolas Bourbaki, 1949– 1973. Notices of the American Mathematical Society, 45(3): 373–380, March 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL https://www.ams.org/
                                        notices/199803/borel.pdf.

Borel:1999:AWA

[Bor99] Armand Borel. André Weil and algebraic topology. Notices of the American Mathematical Society, 46(4):422–427, April 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bott:1991:GR

[Bot91] Raoul Bott. Georges de Rham: 1901–1990. Notices of the American Mathematical Society, 38(2): 114–115, February 1991. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bressoud:1999:HAS

[BP99] David Bressoud and James Propp. How the alternating sign matrix conjecture was solved. Notices of the American Mathematical Society, 46(6):637–646, June/July 1999. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Babai:1998:MPE

[BPV98] László Babai, Carl Pomerance, and Péter Vértesi. The mathematics of Paul Erdős. Notices of the American Mathematical Society, 45(1):19–31, January 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199801/vertesi.
pdf.

Brillhart:1993:DHL

[Bri93] John Brillhart. Derrick Henry Lehmer (1905–1991). Notices of the American Mathematical Society, 40(1):31–32, January 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Brown:1997:CFM

[Bro97] George E. Brown, Jr. Challenges facing mathematics in the twenty-first century. Notices of the American Mathematical Society, 44(5):576–579, May 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199705/
comm-brown.pdf.

Babai:1998:PE

[BS98] László Babai and Joel Spencer.

Paul Erdős (1913–1996). Notices of the American Mathematical Society, 45(1):64–73, January 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199801/
comm-erdos.pdf.

Cartan:1999:AWM

[Car99] Henri Cartan. André Weil: memories of a long friendship. Notices of the American Mathematical Society, 46(6):633–636, June/ July 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Casti:1991:VTP

[Cas91] John L. Casti. VTEX typesetting package: a TEX for tomorrow. Notices of the American Mathematical Society, 38(2):105–109, February 1991. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Casti:1993:B

[Cas93] John L. Casti. Bitmap-free TEX. Notices of the American Mathematical Society, 40(5):464–469, May/June 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). A review of Y&Y’s DVIWindo and DVIP-SONE, and comments on outline fonts for TEX.

Cervone:1995:WTS

[CDK^{+}95] Barbara Tucker Cervone, Bill Duren, J. J. Kohn, J. Laurie Snell, and Marjorie L. Stein. A. W. Tucker: some reminiscences. Notices of the American Mathematical Society, 42 (10):1143–1147, October 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199510/199510-

toc.html. Prepared with the assistance of Alan and Tom Tucker, Based on a memorial resolution written by Harold W. Kuhn.

Choksi:1996:CH

[CFV96] J. R. Choksi, W. H. J. Fuchs, and N. Th. Varopoulos. Carl Herz (1930–1995). Notices of the American Mathematical Society, 43(7):768–771, July 1996. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chandrasekharan:1998:ALS

[Cha98] K. Chandrasekharan. The autobiography of Laurent Schwartz: review of Laurent Schwartz, un mathématicien aux prises avec le siècle [Éditions Odile Jacob, Paris, 1997]. Notices of the American Mathematical Society, 45(9):1141–1147, October 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chern:1996:FGJ

[Che96] Shiing-Shen Chern. Finsler geometry is just Riemannian geometry without the quadratic restriction. Notices of the American Mathematical Society, 43(9): 959–963, September 1996. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chichilnisky:1996:FBM

[Chi96] Graciela Chichilnisky. Fischer Black: the mathematics of uncertainty. Notices of the American Mathematical Society, 43(3):319– 322, March 1996. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Crow:1994:CMK

[CHM94] John Crow, Herb Holden, and Marvin Margolis. Computers and mathematics: Keith Devlin. Notices of the American Mathematical Society, 41(3):195– 201, March 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Church:1992:UWT

[Chu92] P. T. Church. Using workshops to teach TEX. Notices of the American Mathematical Society, 39(8): 842, October 1992. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). AMS-TEX input for two workshop handouts may be obtained by anonymous ftp from syr.edu in the directory /math/papers.

Chern:1998:LA

[CKK^{+}98] S. S. Chern, Thomas Kailath, Bertram Kostant, Calvin C. Moore, and Anna Tsao. Louis Auslander (1928–1997). Notices of the American Mathematical Society, 45(3):390–395, March 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199803/
                                        comm-mem-auslander.pdf.

Christ:1998:APC

[CKSW98] Michael Christ, Carlos E. Kenig, Cora Sadosky, and Guido Weiss. Alberto Pedro Calderón (1920– 1998). Notices of the American Mathematical Society, 45(9): 1148–1153, October 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Coates:1996:AJW

[Coa96] John Coates. [Andrew J.] Wiles receives NAS award in mathematics. Notices of the American Mathematical Society, 43(7):760– 763, July 1996. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Coates:1999:KI

[Coa99] John Coates. Kenkichi Iwasawa (1917–1998). Notices of the American Mathematical Society, 46(10):1221–1225, November 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Coleman:1997:GPD

[Col97] A. John Coleman. Groups and physics — dogmatic opinions of a senior citizen. Notices of the American Mathematical Society, 44(1):8–17, January 1997. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199701/coleman.
pdf.

Constantin:1995:FRO

[Con95] Peter Constantin. A few results and open problems regarding incompressible fluids. Notices of the American Mathematical Society, 42(6):658–663, June 1995.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Crighton:1999:MJL

[CP99] David G. Crighton and T. J. Pedley. Michael James Lighthill (1924–1998). Notices of the American Mathematical Society, 46(10):1226–1229, November 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calderon:1992:AZ

[CS92] A. P. Calderón and Elias M. Stein. Antoni Zygmund: 1900– 1992. Notices of the American Mathematical Society, 39(8):848– 849, October 1992. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Crow:1994:TSR

[CVG94] John Crow, Mario Vassallo, and Gustaf Gripenberg. Three software reviews comprise this month’s column. John Crow presents the second installment of his two-part comparative review of Maple and Macsyma. Mario Vassallo reports on his experiences with spreadsheets and mathematics. Gustaf Gripenberg takes a look at MAX. Notices of the American Mathematical Society, 41(4):299–306, April 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cornish:1998:MSU

[CW98] Neil J. Cornish and Jeffrey R.

Weeks. Measuring the shape of the universe. Notices of the American Mathematical Society, 45(11):1463–1471, December 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199811/
cornish.pdf.

DAmbrosio:1992:ERP

[D’A92] Ubiratan D’Ambrosio. Ethnomathematics: a research program on the history and philosophy of mathematics with pedagogical implications. Notices of the American Mathematical Society, 39(10):1183–1185, November 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Darling:1993:F

[Dar93] R. W. R. Darling. FrameMaker 3.0. Notices of the American Mathematical Society, 40(2):150– 152, February 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Darmon:1999:PFS

[Dar99] Henri Darmon. A proof of the full Shimura–Taniyama–Weil conjecture is announced. Notices of the American Mathematical Society, 46(11):1397–1401, December 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199911/
199911-toc.html.

Daubechies:1995:WAC

[Dau95] Ingrid Daubechies. Wavelets and applications: a celebration of women in mathematics. Notices of the American Mathematical Society, 42 (1):34–35, January 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL http://www.ams.
                                        org/journals/notices/199501/
                                        wim.pdf.

Davis:1997:MM

[Dav97a] Geoff Davis. Mathematicians and the market. Notices of the American Mathematical Society, 44(10):1307–1315, November 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199710/
                                        comm-davis.pdf.

Davis:1997:AIM

[Dav97b] Paul W. Davis. 1996 AMS– IMS–MAA annual survey. Second report. Notices of the American Mathematical Society, 44(8):911–921, September 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199708/
                                        1996-survey.pdf.

Dixmier:1998:LP

[DDH^{+}98] Jacques Dixmier, Michel Duflo, András Hajnal, Richard Kadison, Adam Korányi, Jonathan Rosenberg, and Michèle Vergne. Lajos Pukánszky (1928–1996).

Notices of the American Mathematical Society, 45(4):492–499, April 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199804/
mem-pukanszky.pdf.

Davis:1993:SEM

[DDW93] Joel Davis, Trevian Dray, and André Weideman. The student edition of MATLAB (review). Notices of the American Mathematical Society, 40(4): 329–332, April 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
journals/notices/199304/199304FullIssue.
pdf.

Devlin:1992:AMC

[Dev92a] Keith Devlin. An article by M.

C. Nucci precedes two reviews of the new version 2.0 of the Mathematica system. Notices of the American Mathematical Society, 39(5):427–437, May 1992. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1992:CM

[Dev92b] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 39(8):825–843, October 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1993:CM

[Dev93] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 40(3):236–241, March 1993. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1994:CMa

[Dev94a] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 41(3):195, March 1, 1994. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1994:CMb

[Dev94b] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 41(4):299, April 1, 1994. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1994:CMc

[Dev94c] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 41(5):455, May 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1994:CMd

[Dev94d] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 41(6):611, July 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1994:CMe

[Dev94e] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 41(7):772, September 1994. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Davis:1998:AIM

[DMR98] Paul W. Davis, James W.

Maxwell, and Kinda M. Remick. 1997 AMS–IMS–MAA Annual Survey. (First report). Notices of the American Mathematical Society, 45(1):33–63, January 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199801/
97-ann-surv.pdf.

Davis:1999:ASM

[DMR99] Paul W. Davis, James W.

Maxwell, and Kinda M. Remick. 1998 annual survey of the mathematical sciences. Notices of the American Mathematical Society, 46(8):894–909, September 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199908/
199908-toc.html.

Dolan:1995:BKM

[Dol95]  Louise Dolan.  The beacon

of Kac–Moody symmetry for physics. Notices of the American Mathematical Society, 42 (12):1489–1495, December 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199512/199512-
                                        toc.html.

Doob:1990:TSC

[Doo90] Michael Doob. TEX and the single CPU. Notices of the American Mathematical Society, 37 (3):270–273, March 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dupont:1998:CHS

[DPR^{+}98] Johan Dupont, Anthony Phillips, Vladimir Retakh, Judith Roitman, and Mark Saul. Chih-Han Sah (1934–1997). Notices of the American Mathematical Society, 45(1):79–83, January 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199801/
                                        comm-phillips.pdf.

Dold-Samplonius:1997:IBL

[DS97] Yvonne Dold-Samplonius. Interview with Bartel Leendert van der Waerden. Notices of the American Mathematical Society, 44(3):313–320, March 1997. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With participation of Camilla van der Waerden, Translated from the Italian by Hugo Rossi and Dold-Samplonius.

Dubinsky:1999:RWM

[Dub99] Ed Dubinsky. Review of What is mathematics, really? [Oxford Univ. Press, New York, 1997; MR1462890 (98f:00008)] by R. Hersh. Notices of the American Mathematical Society, 46(9): 1062–1066, October 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Duff:1997:HMC

[Duf97] George F. D. Duff. H. S.

M. Coxeter celebrates 90th birthday. Notices of the American Mathematical Society, 44(3):340–341, March 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199703/
comm-coxeter.pdf.

Duke:1997:SOP

[Duk97] William Duke. Some old problems and new results about quadratic forms. Notices of the American Mathematical Society, 44(2):190–196, February 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1992:CMC

[DW92] K. Devlin and S. X. Wei. Computers and Mathematica column. Notices of the American Mathematical Society, 39(5):428–437, May/June 1992. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Eastwood:1999:VRC

[Eas99] Michael Eastwood. Variations on the de Rham complex. Notices of the American Mathematical Society, 46(11):1368–1376, December 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199911/
199911-toc.html.

Eckmann:1999:TAA

[Eck99] Beno Eckmann. Topology, algebra, analysis — relations and missing links. Notices of the American Mathematical Society, 46(5):520–527, May 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.
ams.org/notices/199905/fea-
eckmann.pdf.

Epstein:1995:EPM

[EL95] David Epstein and Silvio Levy. Experimentation and proof in mathematics. Notices of the American Mathematical Society, 42(6):670–674, June 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Emmer:1997:IEG

[Emm97] Michele Emmer. Interview with Ennio De Giorgi. Notices of the American Mathematical Society, 44(9):1097–1101, October 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199709/

emmer.pdf. Translated from the Italian.

Ewing:1993:MZ

[Ewi93] John Ewing. Max Zorn: 1906– 1993. Notices of the American Mathematical Society, 40(6):640, June/July 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ewing:1996:MCA

[Ewi96] John Ewing. Mathematics: a century ago — a century from now. Notices of the American Mathematical Society, 43(6):663– 672, June 1996. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Faltings:1995:PFL

[Fal95] Gerd Faltings. The proof of Fermat’s last theorem by R. Taylor and A. Wiles. Notices of the American Mathematical Society, 42(7):743–746, July 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Faris:1995:RWL

[Far95] William G. Faris. Random waves and localization. Notices of the American Mathematical Society, 42(8):848–853, August 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Faris:1996:RRO

[Far96a] William Faris. Review of Roland Omnès, The interpretation of quantum mechanics [MR1291602 (95m:81007)]. Notices of the American Mathematical Society, 43(11):1328–1339, November 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Faris:1996:SMS

[Far96b] William Faris. Shadows of the mind: a search for the missing science of consciousness. Notices of the American Mathematical Society, 43(2):203–208, February 1996. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Field:1995:SCH

[FG95] Mike Field and Martin Golubitsky. Symmetric chaos: how and why. Notices of the American Mathematical Society, 42(2): 240–244, February 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199502/199502-
toc.html.

Finn:1999:CSI

[Fin99] Robert Finn. Capillary surface interfaces. Notices of the American Mathematical Society, 46(7): 770–781, August 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fisher:1992:RPB

[Fis92] Stephen D. Fisher. Ralph P. Boas, Jr. (1912–1992). Notices of the American Mathematical Society, 39(10):1189–1190, November 1992. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Friedlander:1999:IFM

[FY99] Susan Friedlander and Victor Yudovich. Instabilities in fluid motion. Notices of the American Mathematical Society, 46 (11):1358–1367, December 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199911/199911-
toc.html.

Gardner:1995:GT

[Gar95] R. J. Gardner. Geometric tomography. Notices of the American Mathematical Society, 42 (4):422–429, April 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199504/199504-
toc.html.

Gillman:1996:MJ

[GH96] Leonard Gillman and Melvin Henriksen. Meyer Jerison (1922– 1995). Notices of the American Mathematical Society, 43(1):17– 18, January 1996. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ghoussoub:1997:PIM

[Gho97] Nassif Ghoussoub. The Pacific Institute for the Mathematical Sciences. Notices of the American Mathematical Society, 44(4):448–450, April 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199704/
comm-pims.pdf.

Golomb:1997:ALW

[GHS97] Solomon Golomb, Theodore Harris, and Jennifer Seberry. Albert Leon Whiteman (1915– 1995). Notices of the American Mathematical Society, 44 (2):217–219, February 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199702/
comm-whiteman.pdf.

Gillman:1997:RC

[Gil97] Leonard Gillman. Rigor in calculus. Notices of the American Mathematical Society, 44 (8):932–934, September 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199708/
comm-gillman.pdf.

Gleason:1990:I

[GJM^{+}90] Andrew M. Gleason, Arthur Jaffe, Barry Mazur, Richard H. Herman, C. Herbert Clemens, János Kollár, Krzysztof Gawȩdzki, Christophe Soulé, and Michael Sipser. ICM-90. Notices of the American Mathematical Society, 37(9):1209–1216, October 1990. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Report on the International Congress of Mathematicians held in Kyoto, August 21– 29, 1990.

Gehring:1998:MLV

[GKO98] Frederick Gehring, Irwin Kra, and Robert Osserman. The mathematics of Lars Valerian Ahlfors. Notices of the American Mathematical Society, 45(2): 233–242, February 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199802/ahlfors.
pdf.

Green:1998:MSR

[GLMM98] Judy Green, Jeanne LaDuke, Saunders Mac Lane, and Uta C. Merzbach. Mina Spiegel Rees (1902–1997). Notices of the American Mathematical Society, 45(7):866–873, August 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gohberg:1990:MGK

[Goh90] Israel Gohberg. Mark Grigorievich Kreĭn, 1907–1989. Notices of the American Mathematical Society, 37(3):284–285, March 1990. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goldreich:1999:P

[Gol99] Oded Goldreich. Pseudorandomness. Notices of the American Mathematical Society, 46 (10):1209–1216, November 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goodman:1998:WSL

[Goo98] Jacob E. Goodman. When is a set of lines in space convex? Notices of the American Mathematical Society, 45(2): 222–232, February 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199802/goodman.
                                        pdf.

Gouvea:1991:LAS

[Gou91] Fernando Q. Gouvêa. A look at PARI and GP. Notices of the American Mathematical Society, 38(8):903–904, October 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Graetzer:1993:A

[Grä93a] George Grätzer. AMS-LaTEX. Notices of the American Mathematical Society, 40(2):148–150, February 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Graetzer:1993:ATI

[Grä93b] George Grätzer. Advances in TEX implementations. I. Post-Script fonts. Notices of the American Mathematical Society, 40(7):834–838, September 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gratzer:1994:ATIa

[Grä94a] George Grätzer. Advances in TEX implementations. II. integrated environments. Notices of the American Mathematical Society, 41(2):106–111, February 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gratzer:1994:ATIb

[Grä94b] George Grätzer. Advances in TEX implementations. III. A new version of LaTEX, finally. Notices of the American Mathematical Society, 41(6):611–615, July/ August 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Graetzer:1994:ATI

[Grä94c] George Grätzer. Advances in TEX: IV. Header and footer control in LaTEX. Notices of the American Mathematical Society, 41(7):772–777, September 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Granville:1997:RBH

[Gra97] Andrew Granville. Review of BBC’s horizon program “Fermat’s Last Theorem”. Notices of the American Mathematical Society, 44(1):26–28, January 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199701/
comm-granville.pdf.

Gromov:1998:PTM

[Gro98] Mikhael Gromov. Possible trends in mathematics in the coming decades. Notices of the American Mathematical Society, 45(7):846– 847, August 1998. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Garfunkel:1990:MOM

[GY90] Solomon A. Garfunkel and Gail S. Young. Mathematics outside of mathematics departments. Notices of the American Mathematical Society, 37(4):408– 411, April 1990. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Garfunkel:1998:SF

[GY98] Solomon A. Garfunkel and Gail S. Young. The sky is falling. Notices of the American Mathematical Society, 45 (2):256–257, February 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199802/
comm-garfunkel.pdf.

Hagedorn:1996:CIB

[Hag96] George A. Hagedorn. Crossing the interface between chemistry and mathematics. Notices of the American Mathematical Society, 43(3):297–299, March 1996. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harrison:1991:ES

[Har91] Jenny Harrison. The Escher staircase. Notices of the American Mathematical Society, 38(7): 730–734, August 1991. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Women in mathematics.

Hartz:1993:M

[Har93] David Hartz. MathType 3.0.

Notices of the American Mathematical Society, 40(5):469–470, May/June 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Review of a Macintosh equation editor front-end to TEX and various word processors.

Helson:1997:MPA

[Hel97] Henry Helson. Mathematics in Poland after the war. Notices of the American Mathematical Society, 44(2):209–212, February 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199702/
comm-helson.pdf.

Henrion:1991:MEL

[Hen91] Claudia Henrion. Merging and emerging lives: Women in mathematics. Notices of the American Mathematical Society, 38(7):724– 729, August 1991. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Women in mathematics.

Hirsch:1998:EHS

[Hir98a] Morris W. Hirsch. Edwin Henry Spanier (1921–1996). Notices of the American Mathematical Society, 45(6):704–705, June/ July 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hirzebruch:1998:KKM

[Hir98b]  Friedrich Hirzebruch.  Kuni-

hiko Kodaira: Mathematician, friend, and teacher. Notices of the American Mathematical Society, 45(11):1456–1462, December 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199811/
hirzebruch.pdf.

Hull:1998:IJA

[HJT98] Thomas Hull, Michael A. Jones, and Diana M. Thomas. Interviewing for a job in academia. Notices of the American Mathematical Society, 45(10):1353– 1357, November 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hoppensteadt:1995:GSM

[Hop95] Frank Hoppensteadt. Getting started in mathematical biology. Notices of the American Mathematical Society, 42(9): 969–975, September 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199509/199509-
toc.html.

Horgan:1994:ADP

[Hor94] John Horgan. Author of “Death of proof” responds to Krantz. Comment on: “The immortality of proof” [Notices Amer. Math. Soc. 41 (1994), no. 1, 10–13; 1 254 635] by S. G. Krantz. Notices of the American Mathematical Society, 41(7):741, August 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Howe:1998:AME

[How98] Roger Howe. The AMS and Mathematics Education: the revision of the NCTM standards. Notices of the American Mathematical Society, 45(2):243– 247, February 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
notices/199802/howe.pdf.

Hwang:1995:WAL

[Hwa95] Andrew D. Hwang. Writing in the age of LaTEX. Notices of the American Mathematical Society, 42(8):878–882, August 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hofer:1991:AF

[HWZ91] Helmut Hofer, Alan Weinstein, and Eduard Zehnder. Andreas Floer: 1956–1991. Notices of the American Mathematical Society, 38(8):910–911, September 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Immerman:1995:DCL

[Imm95] N. Immerman. Descriptive complexity: a logician’s approach to computation. Notices of the American Mathematical Society, 42(10):1127–1133, October 1995. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199510/
199510-toc.html.

Jackson:1990:EWM

[Jac90] Allyn Jackson. Encouraging Women in Math. and Science. Conference focuses on what causes problems and what makes a difference. Notices of the American Mathematical Society, 37(1): 5–6, January 1990. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1991:TPW

[Jac91] Allyn Jackson. Top producers of women mathematics doctorates. Notices of the American Mathematical Society, 38(7):715– 720, August 1991. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Women in mathematics.

Jackson:1992:CFI

[Jac92] Allyn Jackson. Canada’s Fields Institute opens its doors. Notices of the American Mathematical Society, 39(6):564–566, June/ July 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1994:WHN

[Jac94a] Allyn Jackson. What’s Happening: New publication takes the AMS in new direction. Notices of the American Mathematical Society, 41(6):626–627, July/ August 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1994:IZ

[Jac94b] Allyn Jackson. ICM-94 in Zürich.

Notices of the American Mathematical Society, 41(7):763–765, August 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1994:UPF

[Jac94c] Allyn Jackson. Update on proof of Fermat’s Last Theorem. Notices of the American Mathematical Society, 41(3):185–186, March 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jacobowitz:1995:RHC

[Jac95] Howard Jacobowitz. Real hypersurfaces and complex analysis. Notices of the American Mathematical Society, 42(12): 1480–1488, October 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199512/199512-
toc.html.

Jackson:1997:CAO

[Jac97a] Allyn Jackson. Chinese acrobatics, an old-time brewery, and the “Much needed gap”: the life of Mathematical Reviews. Notices of the American Mathematical Society, 44(3):330–337, March 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1997:MWCa

[Jac97b] Allyn Jackson. The Math Wars.

California battles it out over mathematics education reform. I. Notices of the American Mathematical Society, 44(6):695–702, June/July 1997. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1997:MWCb

[Jac97c] Allyn Jackson. The Math Wars. California battles it out over mathematics education reform. II. Notices of the American Mathematical Society, 44(7):817– 823, August 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
notices/199707/comm-calif2.
pdf.

Jackson:1997:MIA

[Jac97d] Allyn Jackson. Mathematics: an interview with AMS President Arthur Jaffe. Notices of the American Mathematical Society, 44(11):1459–1462, December 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1998:BMP

[Jac98a] Allyn Jackson. Bonn’s Max Planck Institute: a new building and a new era. Notices of the American Mathematical Society, 45(5):582–588, May 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199805/planck.
pdf.

Jackson:1998:BGK

[Jac98b] Allyn Jackson. Borcherds, Gowers, Kontsevich, and McMullen receive Fields Medals. Notices of the American Mathematical Society, 45(10):1358–1360, November 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199810/
comm-fields.pdf.

Jackson:1998:DYS

[Jac98c] Allyn Jackson. The demise of the Young Scholars Program. Notices of the American Mathematical Society, 45(3):381–387, March 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199803/
scholars.pdf.

Jackson:1998:PSR

[Jac98d] Allyn Jackson. Peter Shor receives Nevanlinna Prize. Notices of the American Mathematical Society, 45(10):1361, November 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199810/
comm-nevanlinna.pdf.

Jackson:1998:TPM

[Jac98e] Allyn Jackson. Theory into profit: Microsoft invests in mathematicians. Notices of the American Mathematical Society, 45(6): 698–703, June/July 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1999:AIM

[Jac99a] Allyn Jackson. The American Institute of Mathematics. Notices of the American Mathematical Society, 46(5):560–563, May 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1999:DSM

[Jac99b] Allyn Jackson. The dream of a Swedish mathematician: The Mittag-Leffler Institute. Notices of the American Mathematical Society, 46(9):1050–1058, October 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199909/
                                        199909-toc.html.

Jackson:1999:IF

[Jac99c] Allyn Jackson. The IHÉS at forty. Notices of the American Mathematical Society, 46(3): 329–337, March 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL https://www.ams.org/
                                        notices/199903/index.html.

Jackson:1999:IAP

[Jac99d] Allyn Jackson. Interview with AMS president Arthur Jaffe. Notices of the American Mathematical Society, 46(2):221– 223, February 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199902/199902-
                                        toc.html.

Jackson:1999:IHC

[Jac99e] Allyn Jackson. Interview with Henri Cartan [b. 1904]. Notices of the American Mathematical Society, 46(7):782–788, August 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1999:MTS

[Jac99f] Allyn Jackson. Mathematical treasures of the Smithsonian Institution. Notices of the American Mathematical Society, 46 (5):528–534, May 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.
ams.org/notices/199905/fea-
smithsonian.pdf.

Jackson:1999:CMI

[Jac99g] Allyn Jackson. The Clay Mathematics Institute. Notices of the American Mathematical Society, 46(8):888–889, September 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199908/
199908-toc.html.

Jerison:1999:RWA

[Jer99]  David Jerison.  Review of

The world according to wavelets [Peters, Wellesley, MA, 1998; MR1625902 (99g:00029)] by B. B. Hubbard. Notices of the American Mathematical Society, 46(9):1059–1061, October 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1998:ISS

[JK98] Allyn Jackson and Dieter Kotschick. Interview with Shiing Shen Chern. Notices of the American Mathematical Society, 45(7):860– 865, August 1998. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jones:1997:RNG

[JM97] Vaughan Jones and Henri Moscovici. Review of Noncommutative geometry [Academic Press, San Diego, CA, 1994; MR1303779 (95j:46063)] by Alain Connes. Notices of the American Mathematical Society, 44(7):792–799, August 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199707/
                                        jones.pdf.

Jerison:1995:NW

[JS95] David Jerison and Daniel Stroock. Norbert Wiener. Notices of the American Mathematical Society, 42(4):430–438, April 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199504/199504-
                                        toc.html.

Kaplansky:1993:QRD

[Kap93] Irving Kaplansky. Quadratic reciprocity in Dickson’s History. Notices of the American Mathematical Society, 40(9):1155, October 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kass:1996:KM

[Kas96] Seymour Kass. Karl Menger [1902–1985]. Notices of the American Mathematical Society, 43(5):558–561, May 1996. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Keisler:1994:SCK

[Kei94] H. Jerome Keisler. Stephen Cole Kleene 1909–1994. Notices of the American Mathematical Society, 41(7):792–793, August 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kirschner:1996:UMU

[Kir96] Denise Kirschner. Using mathematics to understand HIV immune dynamics. Notices of the American Mathematical Society, 43(2):191–202, February 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Katok:1993:WSM

[KK93] Anatole Katok and Svetlana Katok. Women in Soviet mathematics. Notices of the American Mathematical Society, 40(2): 108–116, February 1993. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Special issue on mathematics in the former Soviet Union.

Knapp:1996:GRHa

[Kna96a] Anthony W. Knapp. Group representations and harmonic analysis from Euler to Langlands. I. Notices of the American Mathematical Society, 43(4):410–415, April 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Knapp:1996:GRHb

[Kna96b] Anthony W. Knapp. Group representations and harmonic analysis. II. Notices of the American Mathematical Society, 43(5):537– 549, May 1996. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Knapp:1999:AWP

[Kna99] Anthony W. Knapp. André Weil: a prologue. Notices of the American Mathematical Society, 46 (4):434–439, April 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Knopp:1990:RPS

[Kno90] Marvin I. Knopp. Rademacher on J(τ), Poincaré series of nonpositive weights and the Eichler cohomology. Notices of the American Mathematical Society, 37(4):385–393, April 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Knuth:1998:LET

[Knu98] Donald E. Knuth. Letter to the editor: Teach calculus with big O. Notices of the American Mathematical Society, 45(6): 687–688, June/July 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Koblitz:1991:MUH

[Kob91] Neal Koblitz. Mathematics under hardship conditions in the Third World. Notices of the American Mathematical Society, 38(9): 1123–1128, October 1991. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kostant:1995:EGT

[Kos95a] Bertram Kostant. Erratum:

“The graph of the truncated icosahedron and the last letter of Galois”. Notices of the American Mathematical Society, 42 (10):1118–1119, October 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See [Kos95b].

Kostant:1995:GTI

[Kos95b] Bertram Kostant. The graph of the truncated icosahedron and the last letter of Galois. Notices of the American Mathematical Society, 42(9):959–968, September 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199509/
199509-toc.html. See erratum

[Kos95a].

Kotschick:1995:GTD

[Kot95] D. Kotschick. Gauge theory is dead! — Long live gauge theory! Notices of the American Mathematical Society, 42(3):335– 338, March 1995. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Klein:1997:CRM

[KR97] David Klein and Jerry Rosen. Calculus reform – for the $ millions. Notices of the American Mathematical Society, 44 (10):1324–1325, November 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199710/
                                        comm-klein.pdf.

Krantz:1994:IPC

[Kra94] Steven G. Krantz. The immortality of proof. Comment on: “The death of proof” [Sci. Amer. 269 (1993), no. 4, 92–103; MR1254647 (95d:00003)] by J. Horgan. Notices of the American Mathematical Society, 41(1):10– 13, January 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See response [Bae94].

Kenig:1998:EBF

[KS98] Carlos E. Kenig and Daniel W. Stroock. Eugene Barry Fabes (1937–1997). Notices of the American Mathematical Society, 45(6):706–708, June/July 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Katz:1999:BD

[KT99] Nicholas M. Katz and John Tate. Bernard Dwork (1923– 1998). Notices of the American Mathematical Society, 46(3):

338–343, March 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
notices/199903/index.html.

Kuperberg:1999:ADS

[Kup99] Krystyna Kuperberg. Aperiodic dynamical systems. Notices of the American Mathematical Society, 46(9):1035–1040, October 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199909/
199909-toc.html.

Lam:1998:RFGa

[Lam98a] T. Y. Lam. Representations of finite groups: a hundred years. I. Notices of the American Mathematical Society, 45 (3):361–372, March 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199803/lam.pdf.

Lam:1998:RFGb

[Lam98b] T. Y. Lam. Representations of finite groups: a hundred years. II. Notices of the American Mathematical Society, 45 (4):465–474, April 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199803/lam.pdf.

Lander:1995:MHUa

[Lan95a] Eric S. Lander. Mapping heredity: using probabilistic models and algorithms to map genes and genomes. Notices of the American Mathematical Society, 42(7):747–753, July 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lander:1995:MHUb

[Lan95b] Eric S. Lander. Mapping heredity: using probabilistic models and algorithms to map genes and genomes. II. Notices of the American Mathematical Society, 42(8):854–858, August 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lang:1995:MRS

[Lan95c] Serge Lang. Mordell’s review, Siegel’s letter to Mordell, Diophantine geometry, and 20th century mathematics. Notices of the American Mathematical Society, 42(3):339–350, March 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lang:1995:SHS

[Lan95d] Serge Lang. Some history of the Shimura–Taniyama conjecture. Notices of the American Mathematical Society, 42(11): 1301–1307, November 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199511/199511-
toc.html.

Lane:1997:PSC

[Lan97a] Neal F. Lane. Problem solving in complex societies. Notices of the American Mathematical Society, 44(5):580–584, May 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199705/
comm-lane.pdf.

Lane:1997:GBS

[Lan97b] Saunders Mac Lane. Garrett Birkhoff (1911–1996) and the survey of modern algebra. Notices of the American Mathematical Society, 44 (11):1438–1439, December 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199711/
comm-maclane.pdf.

Lawson:1990:MK

[Law90] H. Blaine Lawson, Jr. Michio Kuga: 1928–1990. Notices of the American Mathematical Society, 37(5):577, May/June 1990. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lindenstrauss:1994:FMN

[LED^{+}94] Joram Lindenstrauss, Lawrence C.

Evans, Adrien Douady, Aner Shalev, and Nicholas Pippenger. Fields Medals and Nevanlinna Prize presented at ICM-94 in Zürich. Notices of the American Mathematical Society, 41(9): 1103–1111, October 1994. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lehman:1994:PCQ

[Leh94] Eugene Lehman. Pocket calculating the quartic. Notices of the American Mathematical Society, 41(7):777–779, September 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lesniewski:1997:NG

[Les97] Andrew Lesniewski. Noncommutative geometry. Notices of the American Mathematical Society, 44(7):800–805, August 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199707/
                                        lesniewski.pdf.

Lewis:1991:MWU

[Lew91]  D. J. Lewis.  Mathematics

and women: the undergraduate school and pipeline. Notices of the American Mathematical Society, 38(7):721–723, August 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Women in mathematics.

Lepowsky:1999:MWF

[LLMM99] James Lepowsky, Joram Lindenstrauss, Yuri I. Manin, and John Milnor. The mathematical work of the 1998 Fields Medalists. Notices of the American Mathematical Society, 46(1):17–26, January 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Luchins:1996:MOT

[LM96] Edith H. Luchins and Mary Ann McLoughlin. In memoriam: Olga Taussky-Todd [1906–1995]. Notices of the American Mathematical Society, 43(8):838– 847, August 1996. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.
org/journals/notices/199608/
taussky.pdf.

Lions:1997:EG

[LM97] Jacques-Louis Lions and François Murat. Ennio De Giorgi (1928– 1996). Notices of the American Mathematical Society, 44 (9):1095–1096, October 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199709/

murat.pdf. Translated from the French.

Loss:1997:SMI

[Los97] Michael Loss. The stability of matter interacting with fields. Notices of the American Mathematical Society, 44(10):1288– 1293, November 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
notices/199710/loss.pdf.

Lungu:1997:MCS

[LS97] Edward Lungu and Precious Sibanda. Mathematics in the countries of Southern Africa, realities and aspirations. Notices of the American Mathematical Society, 44(5):574–575, May 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199705/
                                        comm-lungu.pdf.

Lieb:1990:VB

[LSW90] Elliott H. Lieb, Elias M. Stein, and Arthur S. Wightman. Valentine Bargmann 1908–1989. Notices of the American Mathematical Society, 37(4):440–441, April 1990. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lui:1997:IVA

[Lui97] S. H. Lui. An interview with Vladimir Arnol’d. Notices of the American Mathematical Society, 44(4):432–438, April 1997. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199704/arnold.
                                        pdf.

Lieb:1998:GES

[LY98] Elliott H. Lieb and Jakob Yngvason. A guide to entropy and the Second Law of Thermodynamics. Notices of the American Mathematical Society, 45(5):571–581, May 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199805/
                                        lieb.pdf.

MacPherson:1993:SRM

[Mac93] Robert MacPherson. Some Russian mathematical institutions.

Notices of the American Mathematical Society, 40(2):119–121, February 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Special issue on mathematics in the former Soviet Union.

MacLane:1995:MGU

[Mac95a] Saunders Mac Lane. Mathematics at Göttingen under the Nazis. Notices of the American Mathematical Society, 42 (10):1134–1138, October 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199510/199510-

toc.html. See correction [Seg96].

Morawetz:1995:RLB

[MAC^{+}95b] Cathleen S. Morawetz, William Abikoff, Carol Corillon, Irwin Kra, Tilla Weinstein, and Jane Gilman. Remembering Lipman Bers. Notices of the American Mathematical Society, 42 (1):8–25, January 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199501/199501-
toc.html.

MacLane:1997:RS

[Mac97a] Saunders Mac Lane. Requiem for the skillful. Notices of the American Mathematical Society, 44(2):207–208, February 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199702/
comm-maclane.pdf.

Mackenzie:1997:FAL

[Mac97b] Dana Mackenzie. Fred Almgren (1933–1997). Lover of mathematics, family, and life’s adventures. Notices of the American Mathematical Society, 44 (9):1102–1106, October 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199709/
                                        comm-almgren.pdf.

MacLane:1997:GBS

[Mac97c] Saunders MacLane. Garrett Birkhoff and the Survey of modern algebra. Notices of the American Mathematical Society, 44 (11):1438–1439, December 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

MacLane:1997:VWM

[Mac97d] Saunders MacLane. Van der Waerden’s Modern Algebra. Notices of the American Mathematical Society, 44(3):321–322, March 1997. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mackenzie:1998:LAA

[Mac98] Dana Mackenzie. Life after academia: How a mathematician found a new career. Notices of the American Mathematical Society, 45(3):401–403, March 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199803/
                                        comm-mackenzie.pdf.

Magid:1999:DGT

[Mag99] Andy R. Magid. Differential Galois theory. Notices of the American Mathematical Society, 46(9): 1041–1049, October 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199909/199909-
toc.html.

Mandrekar:1995:MWN

[Man95] V. Mandrekar. Mathematical work of Norbert Wiener. Notices of the American Mathematical Society, 42(6):664–669, June 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Margolis:1994:MMG

[Mar94] Marvin S. Margolis. MG mathematical graphics system. Notices of the American Mathematical Society, 41(3):200–201, March 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Marker:1996:MTE

[Mar96] David Marker. Model theory and exponentiation. Notices of the American Mathematical Society, 43(7):753–759, July 1996. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mauldin:1997:GFL

[Mau97] R. Daniel Mauldin. A generalization of Fermat’s Last Theorem: the Beal conjecture and prize problem. Notices of the American Mathematical Society, 44(11):1436–1437, December 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199711/
beal.pdf.

Maxwell:1997:CMF

[Max97] James W. Maxwell. Changes in Mathematics faculty composition, fall 1990 to fall 1996. Notices of the American Mathematical Society, 44 (10):1321–1323, November 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199710/
comm-maxwell.pdf.

McDuff:1998:SSN

[McD98] Dusa McDuff. Symplectic structures — a new approach to geometry. Notices of the American Mathematical Society, 45(8): 952–960, September 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199808/mcduff.
pdf.

Milne:1990:FWP

[Mil90] J. S. Milne. Four word processors with TEX capabilities. Notices of the American Mathematical Society, 37(8):1018–1022, October 1990. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Milne:1992:TSW

[Mil92] J. S. Milne. Three scientific word processors. Notices of the American Mathematical Society, 39(8): 838–840, October 1992. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). The word processors Chi-Writer 4.02, Leo 1.4, and Scientific Word 1.0 are discussed in terms of ease of use, effectiveness in producing TEX code, and other features of interest to users who require TEX output but do not wish to use TEX directly.

Milnor:1998:JNB

[Mil98] John Milnor. John Nash, and A beautiful mind [Simon Schuster, New York, 1998; MR1631630 (99j:01021)] by S. Nasar. Notices of the American Mathematical Society, 45(10):1329–1332, November 1998. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Miranda:1999:LSP

[Mir99] Rick Miranda. Linear systems of plane curves. Notices of the American Mathematical Society, 46(2):192–202, February 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199902/199902-
toc.html.

Morgan:1998:RWH

[Mor98] Samuel P. Morgan. Richard Wesley Hamming (1915–1998). Notices of the American Mathematical Society, 45(8):972–977, September 1998. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

                                        http://www.ams.org/notices/
                                        199808/mem-morgan.pdf.

Morawetz:1999:MRR

[Mor99] Cathleen Synge Morawetz. Mathematics to the rescue (retiring Presidential address). Notices of the American Mathematical Society, 46(1):9–16, January 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Minlos:1996:MRD

[MSV96] Robert Minlos, Senya Shlosman, and Nikita Vvedenskaya. Memories of Roland Dobrushin. Translated from the Russian by William Faris. Notices of the American Mathematical Society, 43(4):428–430, April 1996. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Translated from the Russian by William Faris.

McMurran:1999:MC

[MT99] Shawnee McMurran and James Tattersall. Mary Cartwright (1900–1998). Notices of the American Mathematical Society, 46(2):214–220, February 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199902/199902-
                                        toc.html.

Mumford:1997:CRM

[Mum97] David Mumford. Calculus reform — for the millions. Notices of the American Mathematical Society, 44(5):559–563, May 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ni:1998:DCD

[Ni98] Wei-Ming Ni. Diffusion, crossdiffusion, and their spike-layer steady states. Notices of the American Mathematical Society, 45(1):9–18, January 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199801/ni.pdf.

Okerson:1996:WAI

[Oke96] Ann Okerson. Whose article is it anyway? Copyright and intellectual property issues for researchers in the 90s. Notices of the American Mathematical Society, 43(1):8–12, January 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

OShea:1997:DWN

[OP97] Donal O’Shea and Harriet Pollatsek. Do we need prerequisites? Notices of the American Mathematical Society, 44(5):564–570, May 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199705/
199705-toc.html.

Osserman:1999:SPA

[Oss99] Robert Osserman. From Schwarz to Pick to Ahlfors and beyond. Notices of the American Mathematical Society, 46(8):

868–873, September 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199908/199908-
                                        toc.html.

Palais:1999:VMT

[Pal99] Richard S. Palais. The visualization of mathematics: towards a mathematical exploratorium. Notices of the American Mathematical Society, 46(6):647–658, June/July 1999. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Parshall:1996:HWG

[Par96] Karen Hunger Parshall. How we got where we are: an international overview of mathematics in national contexts (1875– 1900). Notices of the American Mathematical Society, 43(3):287– 296, March 1996. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pickover:1991:MBS

[Pic91] C. A. Pickover. Mathematics and beauty: Several short classroom experiments. Notices of the American Mathematical Society, 38(3):192–196, March 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pinsky:1995:PFI

[Pin95] Mark A. Pinsky. Pointwise Fourier inversion in several variables. Notices of the American Mathematical Society, 42(3):330– 334, March 1995. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pomerance:1996:TTS

[Pom96] Carl Pomerance. A tale of two sieves. Notices of the American Mathematical Society, 43 (12):1473–1485, December 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pool:1993:JWG

[Poo93] James C. T. Pool. James Wallace Givens, Jr. (1910–1993). SIAM News, ??(??):??, July 1993. ISSN 0036-1437. Reprinted in [Poo94].

Pool:1994:JWG

[Poo94]  James C. T. Pool.  James

Wallace Givens, Jr. (1910– 1993). Notices of the American Mathematical Society, 41(1): 29–30, January 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
journals/notices/199401/. Reprinted

from [Poo93].

Peskine:1995:MA

[PR95] Christian Peskine and Idun Reiten. Maurice Auslander (1926– 1994). Notices of the American Mathematical Society, 42 (4):450–453, April 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199504/199504-

toc.html. With a note by Hugo Rossi.

Price:1990:CVN

[Pri90] G. Baley Price. Carroll Vincent Newsom: 1904–1990. Notices of the American Mathematical Society, 37(7):877–878, September 1990. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Price:1996:OMO

[Pri96] John F. Price. Optional mathematics is not optional. Notices of the American Mathematical Society, 43(9):964–971, September 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Range:1996:HJB

[RA96] R. Michael Range and Russell W. Anderson. Hans-Joachim Bremermann, 1926–1996. With a biographical sketch of Bremermann by Allyn Jackson. Notices of the American Mathematical Society, 43(9):972–976, September 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With a biographical sketch of Bremermann by Allyn Jackson.

Radin:1995:ST

[Rad95] Charles Radin. Symmetry and tilings. Notices of the American Mathematical Society, 42 (1):26–31, January 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.

org/notices/199501/199501-
toc.html.

Radin:1996:BRM

[Rad96] Charles Radin. Book review of Marjorie Senechal’s Quasicrystals and geometry [Cambridge Univ. Press, Cambridge, 1995; MR1340198 (96c:52038)]. Notices of the American Mathematical Society, 43(4):416–421, April 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raynaud:1999:AWF

[Ray99] Michel Raynaud. André Weil and the foundations of algebraic geometry. Notices of the American Mathematical Society, 46(8): 864–867, September 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199908/199908-
toc.html.

Reid:1996:BJR

[Rei96] Constance Reid. Being Julia Robinson’s sister. Notices of the American Mathematical Society, 43(12):1486–1492, December 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Reiten:1997:CRM

[Rei97] Idun Reiten. Dynkin diagrams and the representation theory of algebras. Notices of the American Mathematical Society, 44 (5):546–556, May 1997. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL https://www.ams.
                                        org/notices/199705/199705-
                                        toc.html.

Ribet:1993:WPT

[Rib93] Kenneth A. Ribet. Wiles proves Taniyama’s conjecture; Fermat’s Last Theorem follows. Notices of the American Mathematical Society, 40(6):575–576, June/ July 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Riley:1998:SME

[Ril98] Richard W. Riley. The state of mathematics education: Building a strong foundation for the 21st century. Notices of the American Mathematical Society, 45(4):487–491, April 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199804/
                                        riley.pdf.

Roberts:1995:PME

[Rob95] Anne Roberts. A partnership in mathematics education. School and university faculty working together. Notices of the American Mathematical Society, 42 (5):547–553, May 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Roitman:1991:WSN

[Roi91] Judith Roitman. What still needs to change (for the good of women in mathematics, and for the good of mathematics). Notices of the American Mathematical Society, 38(7):774–775, August 1991. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Women in mathematics.

Rota:1997:MLL

[Rot97a] Gian-Carlo Rota. The many lives of lattice theory. Notices of the American Mathematical Society, 44(11):1440–1445, December 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199711/
comm-rota.pdf.

Rota:1997:TLW

[Rot97b] Gian-Carlo Rota. Ten lessons I wish I had been taught. Notices of the American Mathematical Society, 44(1):22–25, January 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199701/
comm-rota.pdf.

Rockmore:1999:TPC

[RS99] Daniel Rockmore and J. Laurie Snell. To teach, per chance, to dream. Notices of the American Mathematical Society, 46(8): 874–877, September 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199908/199908-
toc.html.

Rung:1997:SFY

[Run97] Donald C. Rung. A survey of four-year and university mathematics in fall 1995: a hiatus in both enrollment and faculty increases. Notices of the American Mathematical Society, 44(8):923–931, September 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199708/
comm-rung.pdf.

Saari:1995:MCS

[Saa95]  Donald Saari.  Mathemati-

cal complexity of simple economics. Notices of the American Mathematical Society, 42(2): 222–230, February 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199502/199502-
toc.html.

Sadosky:1997:IIE

[Sad97] Cora Sadosky. On issues of immigration and employment for mathematicians. Notices of the American Mathematical Society, 44(11):1474–1477, December 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sarnak:1995:SEC

[Sar95] Peter Sarnak. Selberg’s eigenvalue conjecture. Notices of the American Mathematical Society, 42(11):1272–1277, November 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199511/
199511-toc.html.

Saul:1999:KEH

[Sau99] Mark Saul. Kerosinka: an episode in the history of Soviet mathematics. Notices of the American Mathematical Society, 46(10):1217–1220, November 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shmoys:1996:KSR

[SB96] David B. Shmoys and Steve Batterson. Karp and Smale receive National Medals of Science. Notices of the American Mathematical Society, 43(12):1524–1528, December 1996. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schafer:1991:MWP

[Sch91] Alice T. Schafer. Mathematics and women: Perspectives and progress. Notices of the American Mathematical Society, 38(7): 735–737, August 1991. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Women in mathematics.

Segal:1996:CMG

[Seg96] Sanford L. Segal. Corrections:

“Mathematics at Göttingen under the Nazis” [Notices Amer. Math. Soc. 42 (1995), no. 10, 1134–1138; MR1350011 (96f:01038)] by S. Mac Lane. Notices of the American Mathematical Society, 43(4):405–406, April 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See [Mac95a].

Seress:1997:ICG

[Ser97] Ákos Seress. An introduction to computational group theory. Notices of the American Mathematical Society, 44(6):671–679, June/ July 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199706/
                                        seress.pdf.

Shaker:1992:ACC

[Sha92] Richard J. Shaker. The agency that came in from the cold. Notices of the American Mathematical Society, 39(5):408–411, May 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shimura:1999:AWK

[Shi99] Goro Shimura. André Weil as I knew him. Notices of the American Mathematical Society, 46(4):428–433, April 1999. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Scavo:1993:TM

[SHS93] Tom Scavo, Yannis Haralambous, and Werenfried Spit. TEX on the Macintosh. Notices of the American Mathematical Society, 40(10):1353–1360, December 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Solomon:1995:FSG

[Sol95] Ron Solomon. On finite simple groups and their classification. Notices of the American Mathematical Society, 42(2): 231–239, February 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199502/199502-
toc.html.

Sossinsky:1993:RPM

[Sos93] A. B. Sossinsky. Russian popular math traditions — then and now. Notices of the American Mathematical Society, 40(2):124–128, February 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Special issue on mathematics in the former Soviet Union.

Sossinsky:1997:IUM

[Sos97] A. B. Sossinsky. The Independent University of Moscow in its new quarters. Notices of the American Mathematical Society, 44(4):446–447, April 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199704/
comm-ium.pdf.

Spencer:1998:KK

[Spe98] Donald C. Spencer. Kunihiko Kodaira (1915–1997). Notices of the American Mathematical Society, 45(3):388–389, March 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199803/
comm-obit-spencer.pdf.

Stern:1994:DYR

[ST94] R. Stern and Gang Tian. Donaldson and Yau receive Crafoord Prize. Notices of the American Mathematical Society, 41(7):794– 796, August 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Steinberger:1996:EMJ

[Ste96] Mark Steinberger. Electronic mathematics journals. Notices of the American Mathematical Society, 43(1):13–16, January 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stein:1998:SIR

[Ste98] Elias M. Stein. Singular integrals: the roles of Calderón and Zygmund. Notices of the American Mathematical Society, 45(9): 1130–1140, October 1998. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stoutemyer:1991:CMC

[Sto91] David R. Stoutemyer. Crimes and misdemeanours in the computer algebra trade. Notices of the American Mathematical Society, 38(7):778–785, September 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Strichartz:1999:AF

[Str99] Robert S. Strichartz. Analysis on fractals. Notices of the American Mathematical Society, 46 (10):1199–1208, November 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sumners:1995:LCU

[Sum95] De Witt Sumners. Lifting the curtain: using topology to probe the hidden action of enzymes. Notices of the American Mathematical Society, 42(5):528–537, May 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Singer:1996:TWA

[SW96] I. M. Singer and H. Wu. A tribute to Warren Ambrose. Notices of the American Mathematical Society, 43(4):425–427, April 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saari:1995:IFT

[SX95] Donald G. Saari and Zhihong Xia. Off to infinity in finite time. Notices of the American Mathematical Society, 42(5):538– 546, May 1995. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Teitelbaum:1995:GAS

[Tei95] Jeremy Teitelbaum. The geometry of p-adic symmetric spaces. Notices of the American Mathematical Society, 42(10):1120– 1126, October 1995. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.

org/notices/199510/199510-
toc.html.

Thompson:1997:MSF

[Tho97]  Lisa A. Thompson.  Math-

ematical sciences in the FY 1998 budget. Notices of the American Mathematical Society, 44(7):828–831, August 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199707/
comm-thompson.pdf.

Thomas:1998:UFC

[Tho98] Robin Thomas. An update on the four-color theorem. Notices of the American Mathematical Society, 45(7):848–859, August 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199807/
thomas.pdf.

Toro:1997:DFG

[Tor97] Tatiana Toro. Doubling and flatness: Geometry of measures. Notices of the American Mathematical Society, 44(9):1087–1094, October 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199709/
toro.pdf.

Turner:1991:WRR

[Tur91] Peter R. Turner. Will the “Real” real arithmetic please stand up? Notices of the American Mathematical Society, 38(4):298–304, April 1991. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Taylor:1999:ARH

[TW99] Jean E. Taylor and Sylvia M.

Wiegand. AWM in the 1990s: a recent history of the Association for Women in Mathematics. Notices of the American Mathematical Society, 46(1):27– 38, January 1999. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Uhlenbeck:1995:MSA

[Uhl95] Karen K. Uhlenbeck. Moduli spaces and adiabatic limits. Notices of the American Mathematical Society, 42(1):41–42, January 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

VanderWaerden:1997:MMW

[Van97] B. L. Van der Waerden. From matrix mechanics and wave mechanics to unified quantum mechanics. Notices of the American Mathematical Society, 44(3):323–329, March 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199703/
vanderwaerden.pdf.

Varadarajan:1998:NVP

[Var98] V. S. Varadarajan. Niels Vigand Pedersen (1949–1996). Notices of the American Mathematical Society, 45(1):74–78, January 1998. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199801/
comm-pedersen.pdf.

Varadarajan:1999:ERA

[Var99]  V. S. Varadarajan.  Essay

review of André Weil: The apprenticeship of a mathematician [Birkhäuser, Basel, 1992; MR1139519 (93j:01015b)]. Notices of the American Mathematical Society, 46(4):448–456, April 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Veldkamp:1991:HF

[Vel91] Ferdinand D. Veldkamp. Hans Freudenthal: 1905–1990. Notices of the American Mathematical Society, 38(2):113–114, February 1991. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vitulli:1997:WGA

[VF97] Marie A. Vitulli and Mary E. Flahive. Are women getting all the jobs? Notices of the American Mathematical Society, 44(3):338–339, March 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199703/
comm-vitulli.pdf.

Walker:1997:FAW

[Wal97] James S. Walker. Fourier analysis and wavelet analysis. Notices of the American Mathematical Society, 44(6):658–670, June/ July 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199706/
walker.pdf.

Wayne:1997:PSN

[Way97] C. Eugene Wayne. Periodic solutions of nonlinear partial differential equations. Notices of the American Mathematical Society, 44(8):895–902, September 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199708/
wayne.pdf.

Wilson:1996:WLC

[WCA^{+}96] W. Stephen Wilson, S. S. Chern, Shreeram S. Abhyankar, Serge Lang, and Jun ichi Igusa. Wei-Liang Chow [1911–1995]. Notices of the American Mathematical Society, 43(10):1117–1124, October 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Weibel:1996:RWT

[Wei96a] Charles A. Weibel. Robert W.

Thomason (1952–1995). Notices of the American Mathematical Society, 43(8):860–862, August 1996. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Weinstein:1996:GUI

[Wei96b] Alan Weinstein. Groupoids: unifying internal and external symmetry. A tour through some examples. Notices of the American Mathematical Society, 43(7):744– 752, July 1996. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

White:1997:MFJ

[Whi97] Brian White. The mathematics of F. J. Almgren, Jr. Notices of the American Mathematical Society, 44(11):1451–1456, December 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199711/
                                        comm-white.pdf.

Wightman:1995:EPW

[Wig95] A. S. Wightman. Eugene Paul Wigner (1902–1995). Notices of the American Mathematical Society, 42(7):769–771, July 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Williams:1997:DS

[Wil97] H. C. Williams. Daniel Shanks (1917–1996). Notices of the American Mathematical Society, 44(7):813–816, August 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/199707/
                                        comm-shanks.pdf.

Witten:1998:MMM

[Wit98] Edward Witten. Magic, mystery, and matrix. Notices of the American Mathematical Society, 45(9):1124–1129, October 1998. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199809/
witten.pdf.

Woolf:1990:MRD

[Woo90] William B. Woolf. The Mathematical Reviews database: the power of modern technology. Notices of the American Mathematical Society, 37(2):132–134, February 1990. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Willinger:1998:WMM

[WP98] Walter Willinger and Vern Paxson. Where mathematics meets the Internet. Notices of the American Mathematical Society, 45(8):961–970, September 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/notices/199808/paxson.
pdf.

Yorke:1997:EMC

[YH97] J. A. Yorke and M. D. Hartl. Efficient methods for covering material and keys to infinity. Notices of the American Mathematical Society, 44(6):685–687, June/ July 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Young:1997:GBA

[You97a] David M. Young. Garrett Birkhoff and applied mathematics. Notices of the American Mathematical Society, 44(11):

1446–1450, December 1997. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.ams.
org/journals/notices/199712/
199712FullIssue.pdf;   https:
//www.ams.org/notices/199711/
comm-young.pdf.

Youngen:1997:TMM

[You97b] Ralph Youngen. Toward a mathematical markup language. Notices of the American Mathematical Society, 44(9):1107–1109, October 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199709/
comm-youngen.pdf.

Young:1998:DCD

[You98] Lai-Sang Young. Developments in chaotic dynamics. Notices of the American Mathematical Society, 45(10):1318–1328, November 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199810/
young.pdf.

Zassenhaus:1990:MHJ

[Zas90] Hans Zassenhaus. Marshall Hall, Jr.: 1910–1990. Notices of the American Mathematical Society, 37(8):1033, September 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zeilberger:1993:TPT

[Zei93] Doron Zeilberger. Theorems for a price: Tomorrow’s semi-rigorous mathematical culture. Notices of the American Mathematical Society, 40(8):978–981, September 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zworski:1999:RPG

[Zwo99] Maciej Zworski. Resonances in physics and geometry. Notices of the American Mathematical Society, 46(3):319–328, March 1999. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/199903/
index.html.